### v1

# MMSearch-Plus: A Simple Yet Challenging Benchmark for Multimodal Browsing Agents

Abstract
Large multimodal language models (MLLMs) are increasingly deployed as web agents, yet many multimodal browsing benchmarks can be solved by shallow, fixed workflows that lean on high-recall image search and nearby text—masking the genuinely multimodal challenges of fine-grained visual reasoning, provenance verification, and long-horizon tool use. We introduce MMSearch-Plus, a benchmark of 311 tasks that highly demands multimodal understanding while preserving the difficulty profile of strong text-only browsing suites. Each item is constructed to contain multiple weak, localized visual signals that must be extracted, propagated through iterative text–image search, and cross-validated under retrieval noise before answering. Our curation procedure, Spatial–Temporal Extrapolation, seeds questions whose answers require extrapolating from spatial cues (micro-text, part-level appearance, layouts, signage) and temporal traces (broadcast overlays, seasonal context) to out-of-image facts such as events, dates, and venues. We provide a model-agnostic agent framework with browsing tools, and evaluate a range of closed and open MLLMs. The strongest agent (o3) attains 15.1% without search and 36.0% accuracy with rollout under our framework, while a strong open-source model (Qwen-2.5-VL-72B-Instruct) achieves 0.0% without search and 6.9% after 20 rounds of search. Beyond answer accuracy, we assess bounding-box production and cropped-image search, and conduct an error analysis that surfaces failures in source verification, part-based reasoning, and long-horizon planning.
Project Page: https://mmsearch-plus.github.io
footnotetext: *Equal contribution. Corresponding authors: liu.rui2@huawei.com, lpk@cs.hku.hk.
1 Introduction
Figure 1: Example MMSearch-Plus item. Given a 2025 concert photo and the query “What was the singer’s performance time?”, the agent’s golden path extracts localized cues—micro-text/lyrics, a crop of the singer, festival/brand signage, and a distinctive stage prop—and issues targeted image/text searches that (i) identify the artist/outfit, (ii) resolve the event (e.g., a specific festival and city/date), and (iii) retrieve and cross-validate the official schedule to obtain the exact performance time. The answer is not explicitly written in the prompt or the image, underscoring MMSearch-Plus’s emphasis on fine-grained multimodal reasoning with provenance checks under retrieval noise.
Recent advances in large multimodal language models (MLLMs) have enabled them to act as capable agents that integrate language, vision, and tool use to accomplish complex tasks. A growing body of work studies how MLLMs leverage external search engines—both text- and image-based—to answer information-seeking questions.
Multimodal browsing benchmarks such as MMSearch (Jiang et al., 2024) primarily test a model’s ability to coordinate text and images. However, many MMSearch-style questions can be solved by relatively fixed workflows that require little genuine multimodal reasoning. In particular, MMSearch heavily relies on external image search: the MLLM often plays the role of an orchestrator rather than a visual reasoner. When the search engine retrieves highly relevant images, even a unimodal LLM—without visual perception—can frequently answer by reasoning over the accompanying page text alone. This occurs because a single strong image search can surface pages whose surrounding text already contains the answer. In such cases, the image search tool and the MLLM are partially interchangeable as sources of image-related information; the only distinctly multimodal step is often image–source cross-validation, a narrow slice of multimodal reasoning.
Empirically, MMSearch requires short reasoning chains and achieves relatively high model scores: GPT-4o exceeds 60% on MMSearch (end-to-end VQA), and MMSearch-R1-7B reports 54.6% average accuracy (Wu et al., 2025). By contrast, recent text-only browsing benchmarks emphasize persistence and creative, multi-step search. BrowseComp (Wei et al., 2025) targets web navigation for hard-to-find, entangled information; on this task, GPT-4o and GPT-4.5 each score below 1% in direct-answer settings and under 2% even with browsing tools (Wei et al., 2025). “Deep Research” reports 51.5% on BrowseComp, though it was trained for the benchmark and public implementation details remain limited, while open-source agents such as WebSailor reach roughly 12% after substantial engineering (Li et al., 2025a). Taken together, these results suggest that current multimodal benchmarks may not yet match the difficulty of complex text-only browsing, motivating a BrowseComp-style multimodal benchmark with similarly demanding reasoning.
What makes a challenging benchmark for multimodal browsing agents.
Building on these insights, we argue that a strong multimodal benchmark should: (i) require fine-grained, exhaustive visual reasoning that compels models to mine subtle, localized cues rather than rely on a single salient (“dominant”) entity that occupies most of the image; and (ii) enforce provenance and source verification under retrieval noise—i.e., discriminating authentic sources when image results are noisy or conflicting and validating images embedded in webpages. Concretely, the benchmark should stress robust cross-modal evidence gathering, careful verification, and resilience to near-duplicates. For example, two football matches from adjacent seasons may share nearly identical visuals (kits, stadium angles). Correctly answering requires inspecting localized markers (e.g., patch details, sponsor variants, scoreboard typography) and cross-checking dates and venues against trustworthy sources. In real applications, difficult multimodal information-seeking tasks exhibit three recurring challenges that our benchmark targets:
  1. 1.
    Noisy or conflicting retrieval. Image search may return non-relevant or contradictory results; agents must identify authentic sources via fine-grained multimodal understanding and cross-source consistency checks.
  2. 2.
    Exhaustive, part-based visual reasoning. When holistic matching fails, agents must reason over parts and subregions—often iteratively cropping and re-searching—to piece together fragmented clues that identify the target entity or event.
  3. 3.
    Professional multimodal handling and long, tool-augmented chains. Agents must extract answers from heterogeneous media (images, documents, videos) across the web, sustain long reasoning/tool-use chains, and invoke visual tools programmatically (e.g., cropping, OCR) when helpful, akin to advanced agents such as o3/GPT-5-Thinking.
Our benchmark.
Our work, MMSearch-Plus, aims to bridge the gap by combining the persistence and high-reasoning demands of BrowseComp with truly multimodal workflows. Unlike (Geng et al., 2025), where images often collapse to a single identifiable entity, images in our benchmark contain multiple subtle visual signals that must be connected throughout the entire reasoning process. To systematically curate hard multimodal search questions, we introduce Spatial-Temporal Extrapolation, a principled curation procedure that seeds questions whose answers require extrapolating from spatial (localized visual cues, layouts, micro-text, uniforms, signage) and temporal signals (broadcast overlays, seasonal context, historical rosters/schedules) to identify events, dates, or locations not explicitly present in the prompt or image. This forces models to extract fine-grained cues, propagate them through iterative (image/text) searches, and verify provenance amidst retrieval noise—directly addressing the challenges above.
MMSearch-Plus targets realistic scenarios where answers cannot be read off from the prompt or the image alone. For example, given a screenshot from an esports broadcast, the agent may need to infer the match date—absent from both the query and the image—by integrating subtle overlays with external evidence. Similar patterns arise in sports, geolocation, and lifestyle vlogs, domains we explicitly include to move beyond typical web imagery.
1.1 Contributions
Our contributions are as follows:
  • We propose MMSearch-Plus, a challenging benchmark that integrates fine-grained multimodal reasoning (with sparse, localized visual signals) and external tool use, while maintaining the difficulty profile of text-based browsing tasks like BrowseComp (Wei et al., 2025).
  • We introduce the curation procedure Spatial-Temporal Extrapolation to construct difficult multimodal search questions that require extrapolating from localized spatial cues and temporal signals to out-of-image facts (e.g., events, dates, venues), thereby stressing provenance verification under retrieval noise.
  • We develop a custom search agent framework and provide a comprehensive evaluation of current MLLMs in this setting, revealing gaps and opportunities for agents capable of sophisticated, tool-augmented multimodal reasoning.
A concurrent work, BrowseComp-VL (Geng et al., 2025), increases difficulty mainly by expanding the text search space in a BrowseComp-like manner, while the image often reduces to a single identifiable entity that is straightforwardly searchable and quickly disposable. More concretely, BrowseComp-VL from (Geng et al., 2025) is built by first creating multi-hop text QA (BrowseComp-style, with entity obfuscation) and then converting to VQA by replacing the explicit entity mention with an image of that entity (via web image retrieval). As a result, many problems primarily hinge on text search and webpage visiting after an initial visual anchoring step, rather than sustained fine-grained visual reasoning. Another concurrent work,  (Li et al., 2025b), is related. Our contributions differ in (a) data sources; (b) a general search framework that supports any MLLM; and (c) an analysis of whether thinking with images/cropping helps current MLLMs excel on our benchmark.
2 Method
2.1 Spatial-Temporal Extrapolation
A central challenge in BrowseComp-like benchmarks arises from the large intermediate search space induced by soft, fuzzy constraints. This requires agents to perform non-trivial cross-validation and identify the correct target. In designing our benchmark, rather than remixing existing text-only datasets, we aim to construct problems that naturally expand the search space during multimodal information seeking, thereby testing an agent’s ability for strategic planning and uncertainty-aware reasoning in dynamic environments.
Inspired by Geoguessr-style tasks, our problems are anchored on real-world events. Agents must piece together fragmented visual information to identify the underlying source event. Task difficulty is modulated by varying the richness of both visual clues and textual context. Even a single visual fragment can expand the search space dramatically, requiring careful comparison with retrieved content and cross-validation against other multimodal evidence. In more difficult cases, this mirrors human cognition: the agent must iteratively generate hypotheses, verify them against internal knowledge or retrieved content, and refine its reasoning chain across interleaved text and images. Such processes result in extended trajectories that demand robust contextual understanding.
Figure 2: Examples of spatial and temporal relations for question curation.
Once an event is identified, we formulate questions that probe its metadata or chain together multi-hop queries. To further elevate difficulty, we introduce Spatial-Temporal Extrapolation. Instead of asking what is directly visible, we query what is contextually implied but physically absent, compelling reasoning beyond the pixels to reconstruct the broader event. Spatial extrapolation targets unseen entities—individuals off-screen, facing away, or partially obscured—while temporal extrapolation probes events preceding or following the depicted moment. This design forces agents to first localize the event precisely (e.g., time, match, or episode), and then retrieve and reason over wider contextual knowledge from diverse sources.
2.2 Data Collection
Building on the Spatial-Temporal Extrapolation framework, we instantiated the data collection process as follows.
We began with a set of predefined categories and generated search keywords using either LLM assistance or manual curation. Data were primarily collected from three sources: YouTube, Bilibili, and ArXiv. Keywords from categories such as Vlog were searched on YouTube and Bilibili, while those in the Academic category were searched on ArXiv, where relevant research papers (PDFs) were retrieved. We treat both videos and papers as events.
For videos, human annotators manually selected keyframes suitable for constructing questions. The keyframes underwent deanoymouization for sensitive content. For papers, annotators extracted figures or tables, capturing them via screenshots. A screenshot was considered suitable if it (i) contained vague or noisy information that challenged perception, or (ii) depicted less familiar entities requiring external retrieval to answer associated questions. This ensured that events were either (a) novel, i.e., published after model knowledge cutoffs, or (b) rarely represented in training corpora, consistent with the principles in (Jiang et al., 2024).
Filtering.
Although the collection focused on new or rarely seen events, we observed that certain closed-source models (e.g., GPT-4o, GPT-5, Gemini-2.5-Pro) could occasionally solve questions without external search. To counteract this, we employed an adversarial filtering procedure to increase difficulty. Specifically, annotators removed or obfuscated questions answerable without retrieval through the following methods:
  • Cross-validation: Annotators designed problems outside their own knowledge base, then verified solvability by testing on at least two closed-source MLLMs.
  • Image perturbation: Critical visual tokens (e.g., unique logos, labels) were blurred or masked.
  • Text sanitization: Explicit metadata (e.g., titles, author names) was redacted while maintaining semantic content.
  • Iterative refinement: Questions that remained trivially solvable were discarded.
2.3 Dataset Statistics
Metric Value
#Total Entries 311
#Total Entries w/ Images 311
Avg #Images per Entry 1.418
Avg #Words in Question 12.273
Avg #Words in Answer 3.840
#Primary Categories 8
#Secondary Categories 43
Figure 3: (a) Distribution of primary categories in MMSearch-Plus. (b) Summary of statistics.
Figure 3 depicts the distribution of MMSearch-Plus tasks across eight primary categories, which is approximately balanced. Each task includes at least one image, with additional images provided when a single view is insufficient. The benchmark primarily comprises short-answer questions, with answers averaging 3.8 words. In hindsight, we labeled the difficulty level of a task as Easy if it can be answered correctly without search or with only image search by o3 or Gemini. Otherwise, the task is labeled as Difficult (interchangeably Hard).
2.4 Benchmark Metric and Evaluation
We adopt accuracy as the primary evaluation metric. For each question, a set of acceptable answers is provided. We use LLM-as-a-judge to compare model predictions against this set, deeming an answer correct if it matches any acceptable variant. Validation through manual inspection showed complete agreement between human judgment and LLM assessment.
We initially tested both GPT-4o and Gemini-2.5-Pro as judges and found their outputs consistent. The subsequent experiments used GPT-4o. To further support reproducibility, we optionally provide lightweight rule-based verification methods.
2.5 Search Agent Framework
We leveraged SerpApi to support both text and image search. The framework is compatible with sub-image cropping and sub-image search for fine-grained analysis. However, due to the limited grounding ability of the tested MLLMs, we did not incorporate sub-image cropping in the main experiments.
The framework incorporated several optimizations:
  • Search result caching: For each image in MMSearch-Plus, we cache the corresponding image-search results (i.e., the ranked URLs and thumbnails returned by the engine). This reduces latency and token cost during repeated agent steps and evaluation, and it does not deterioriate reproducibility because the top image-search results are relatively stable across time in our setting.
  • Top-$K$ search results retrieval: The most relevant search results determined by the search engine are passed to the MLLMs.
  • Content extraction: Retrieved webpages are summarized by an MLLM into web_info and related_info, discarding boilerplate HTML to reduce token cost. web_info contains information relevant to the question, whereas related_info captures the relevance between a webpage’s thumbnail and the query image.
  • Contextualized evaluation: The agent is always provided with the full conversation history and prior reasoning steps.
Table 1: End-to-end results on the MMSearch+ benchmark, across search modes. All numbers are accuracy (%).
Model / Search Mode Avg By Category Difficulty
Geo. Sports Acad. Film/TV Tech Games Vlog Music Easy Hard
Closed-source LMMs
o3 (2025-04-16)
Without Search 15.1 31.2 14.8 6.0 17.5 13.9 3.2 5.3 11.8 50.0 0.0
Image Search 19.3 28.1 14.8 18.0 30.0 22.2 3.2 5.3 17.6 63.8 0.0
Full Rollout 36.0 35.9 24.1 50.0 42.5 44.4 16.1 42.1 29.4 54.3 28.1
GPT-5
Without Search 10.3 21.9 7.4 4.0 7.5 8.3 0.0 5.3 15.8 27.7 2.8
Image Search 16.4 25.0 11.1 14.0 22.5 19.4 3.2 0.0 29.4 50.0 1.8
Gemini-2.5-Pro
Without Search 10.6 15.6 11.1 6.0 12.5 13.9 0.0 15.8 5.9 35.1 0.0
Image Search 16.4 26.6 11.1 18.0 20.0 16.7 3.2 0.0 23.5 54.3 0.0
Full Rollout 23.8 39.1 14.8 12.0 27.5 33.3 6.5 26.3 29.4 46.8 13.8
Open-source LMMs
Qwen-2.5-VL-72B-Instruct
Without Search 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Image Search 13.5 20.3 7.4 18.0 17.5 11.1 3.2 0.0 23.5 41.5 1.4
Full Rollout 6.1 9.5 7.4 4.0 5.0 2.8 3.2 5.3 11.8 17.0 1.4
3 Experiment
3.1 Experiment Setup
Evaluated Models.
We evaluate both closed-source and open-source MLLMs on our benchmark. Closed-source models include Gemini-2.5-Pro (hereafter Gemini), o3, and GPT-5, representing the current frontier of proprietary multimodal systems. On the open-source side, we focus on Qwen-2.5-VL-72B-Instruct (hereafter Qwen), a state-of-the-art vision–language model trained at large scale. All models are used in their publicly released inference versions without any task-specific fine-tuning.
Implementation Details.
To ensure a fair comparison, we standardize the search-and-reasoning pipeline across all models. We consider three search modes:
  1. (a)
    Without Search. The MLLM answers using only the question and image inputs.
  2. (b)
    Image Search. For each benchmark image, the top-10 image-search results are summarized by Gemini and cached. The MLLM then receives the question, the images, and these summaries as input. For multi-image questions, we append each image’s summaries to the model context.
  3. (c)
    Full Rollout. The model runs our search-agent framework and may call search tools as needed. Image-search settings match those in Image Search. For text search, we retrieve the top-5 websites by default. The process is capped at 20 search rounds (combining text and image search) to balance efficiency and coverage. In each text-search round, the model generates 3–5 refined queries; in each image-search round, it issues exactly one query. For Qwen, pilot studies showed that more than 10 rounds yield no further gains, so we cap its search at 10 rounds.
4 Results
The results of MLLMs across search modes are presented in Table 1.
Baseline Performance (Without Search)
Without any external retrieval, the open-source model (Qwen) achieved 0.0% across all categories, indicating that the benchmark is intentionally adversarial to parametric-only knowledge. Closed-source baselines fare better but are still far from saturation: GPT-5 at 10.3%, Gemini-2.5-Pro at 10.6%, and o3 at 15.1%. The category breakdown reveals why: even for the strongest no-search model (o3), questions that likely require long-tail facts or fine-grained entity disambiguation remain difficult (e.g., Games 3.2, Vlog 5.3, Acad. 6.0), whereas more visually grounded or canonical domains such as Geo. (31.2) and the Easy split (50.0) are comparatively tractable. This spread suggests that, on MMSearch+, in-context parametric knowledge alone is insufficient: many items appear to require explicit grounding to web evidence, especially where named entities, screen text, or episodic, pop-culture–style facts dominate.
Figure 4: Accuracy of o3 and Gemini-2.5-Pro across 3 search modes.
Effect of Image Search
A single round of image search provides a consistent, modest lift: on average +4.2 points for o3 (from 15.1 to 19.3), +6.1 for GPT-5 (10.3 to 16.4), and +5.8 for Gemini-2.5-Pro (10.6 to 16.4). Qwen also benefits, jumping from 0.0 to 13.5 (+13.5). Overall, the pattern implies that a single, reasonably relevant image can unlock coarse disambiguation for Qwen but leaves complex, compositional cases unresolved.
Full Rollout
Allowing iterative search yields a step-change in performance. o3 reaches 36.0% (absolute +20.9 over no-search), and Gemini-2.5-Pro reaches 23.8% (+13.2). The pattern aligns with tasks that require progressive narrowing: o3’s largest category deltas come in Acad. (+44.0; 6.0 $\rightarrow$ 50.0), Vlog (+36.8), and Tech (+30.5), where multi-hop browsing can surface specific frames, quotes, or specs. Gemini also shows substantial lifts in Geo. and Music (both +23.5), Tech (+19.4), and Film/TV (+15.0). By contrast, Qwen regresses under our rollout controller: its average drops from 13.5 (image search) to 6.1 (–7.4). This suggests that iterative search is not plug-and-play for Qwen: without model-specific tuning of the rollout policy (e.g., hop budgeting, snippet ranking, stopping criteria) and/or stronger web-reading robustness, extra hops can introduce distractors faster than they add evidence.
5 Analysis
5.1 Rollout Trajectory Statistics
Figure 5: Reasoning-trajectory statistics. Rows: MLLMs (o3, Gemini, Qwen from top to down). Columns: left—distribution of image search calls and text search calls per trajectory; right—number of words in an MLLM’s response versus the number of search calls, stratified by correctness. Statistics collected in full rollout mode.
Figure 5 provides an overview of reasoning-trajectory statistics across three MLLMs (o3, Gemini, Qwen): the left column reports the distribution of text- and image-search calls per conversation, while the right column relates assistant word count to the number of search calls, stratified by correctness. All statistics are collected in full rollout mode. From the figure, we summarize:
  • More calls when wrong. Incorrect trajectories typically include more text-search calls than correct ones, indicating that uncertainty triggers extra querying without surfacing decisive evidence.
  • Gemini under-explores. Despite a budget of $20$ calls, Gemini usually issues fewer than $5$ search calls per trajectory; text-search medians hover near 011 and image-search is near zero for both correct and incorrect cases. When it does search, response length grows mildly with call count, suggesting reliance on parametric knowledge rather than iterative retrieval.
  • Little link between calls and verbosity. Across models there is no robust monotonic relationship between the number of calls and the length of the assistant’s response. Trends are model-dependent: o3 shows a slight negative slope, Gemini a mild positive slope, and Qwen’s incorrect trajectories become more verbose with more calls while correct ones do not.
5.2 Observations
In addition to the result analysis from Section 4 that focuses on search modes and difficulties, we highlight three themes that relates to world knowledge, context window and visual clue grounding.
Closed-source MLLMs exhibit broad world knowledge, especially geographic.
When prompted with rare or niche events (e.g., images from small or sparsely documented locales), some closed-source models correctly recognize the event and provide relevant context, indicating substantial factual knowledge beyond mainstream datasets. By category, the geography tasks achieve the highest accuracies, consistent with findings from GeoBench (ccmdi, 2023).
Limitations remain in grounded reasoning.
Despite memory advantages, models often struggle with spatial grounding—e.g., producing correct bounding boxes for downstream search. Enabling agent-driven cropping frequently deteriorates performance on our benchmark. Consequently, we did not run a full-scale experiment with an image-cropping tool.
o3 shows strong memory retention.
Across several examples, o3 retains context over long interaction windows—successfully coordinating over 10+10{+} search rounds (Figure 5) and considering $10\times 5$ (top-5) $\approx 50{+}$ retrieved items. This persistence improves factual consistency and reduces redundant tool calls, leading to smoother task completion in information-seeking settings.
Qwen sometimes cannot call tools properly.
The bottom row of Figure 5 shows a long-tailed distribution of Qwen’s image-search calls—occasionally spiking to many calls within a single trajectory—while text-search calls remain low. On the right, incorrect trajectories exhibit increasing verbosity with more calls, whereas correct ones do not benefit from additional calls and sometimes even shorten, reflecting unstable tool-use dynamics. In practice, we also observe execution failures: out of $311$ tasks in MMSearch-Plus, Qwen produced invalid tool calls on $45$ tasks, for a total of $421$ invalid calls. These failures likely trigger cascaded retries (inflating call counts) without improving evidence quality, contributing to the observed variance and degraded accuracy.
Figure 6: Representative error case studies. Each panel corresponds to a high-frequency category in our taxonomy (Figure 7). Panels include the prompt, ground-truth event/answer, assigned error type, and a shortened rollout. (a) Key information not extracted: the summarizer omits required facts from retrieved pages. (b) Relevance not verified: a visually plausible page is retrieved, but the predicted winner corresponds to the wrong event (e.g., 2021 vs. 2024 snooker match). (c) Question misunderstood: the browsing agent misreads multimodal context and answers prematurely. (d) May require video understanding: the required datum exists only in narration from the source video, but our current search agent framework does not require video input.
5.3 Error Analysis
We categorize failure cases using a manually curated taxonomy and report the distribution over categories in Figure 7. Labels are assigned by the authors (single dominant label per erroneous prediction) for Gemini-2.5-Pro under the full-rollout setting.
High-frequency categories.
The two most prevalent failure modes are no relevant information found and hallucination, responsible for $51.1\%$ and $11.5\%$ of errors, respectively ($62.6\%$ combined). The former typically reflects a retrieval failure: noisy or overly specific queries miss on-topic pages. The latter reflects vision-driven misattribution: the model latches onto an incorrect event from internal knowledge and either does not invoke search or discounts contradictory evidence in the retrieved pages. Figure 6 provides examples of four additional high-frequency types.
Other observed categories.
We also observe failures from misinterpreting the multimodal context; hallucinations grounded in either images or parametric knowledge; refusals or multi-answer outputs when evidence is insufficient to disambiguate a single event; failing to invoke search (falling back on outdated priors); and errors when summarizing relevant webpages.
Figure 7: Distribution of error types. Manually annotated on Gemini-2.5-Pro (full rollout).
Qualitative takeaways and implications.
Two recurring patterns emerge: (i) there is no single “golden path”—agents can succeed via varied orders of search/crop/verify; and (ii) stronger multimodal internal knowledge narrows the practical search space, enabling targeted verification over broad exploration. Accordingly, evaluation and training should reward families of effective strategies rather than a single script. While augmenting multimodal internal knowledge can reduce exploration overhead, end-to-end reliability still depends on robust verification and tighter cross-modal grounding.
6 Related Work
Multimodal Reasoning Benchmarks.
Classical vision-and-language benchmarks such as CLEVR (Johnson et al., 2017), VCR (Zellers et al., 2019), Visual Genome (Krishna et al., 2016), and InfographicVQA (Mathew et al., 2021) established core evaluations for compositional reasoning, commonsense inference, and visual question answering. Recent efforts extend these settings to longer contexts and interleaved modalities, including LongVideoBench (Wu et al., 2024), MVBench (Li et al., 2024), Video-MMMU (Hu et al., 2025), and VideoRAG (Jeong et al., 2025), which highlight long-context understanding and retrieval augmentation. Other work such as MIRB (Zhao et al., 2024), pixel-level VCR (Zhang et al., 2024), and ReasonMap (Feng et al., 2025) aim to connect perception with explicit reasoning. However, these datasets are primarily self-contained and do not require dynamic interaction with the open web or tool-mediated search.
Web Browsing and Search-Enhanced Agents.
A large body of work explores language models as web agents that plan, browse, and gather evidence. Earilier works like FRAMES (Krishna et al., 2024) focus on multi-hop QAs on Wikipedia-like data. More recently, benchmarks such as BrowseComp (Wei et al., 2025) and its Chinese variant BrowseComp-ZH (Zhou et al., 2025) test multi-step persistence in text-heavy web navigation, where even frontier models struggle. Open-source agents including WebSailor (Li et al., 2025a), WebShaper (Tao et al., 2025), and WebWatcher (BrowseComp-VL) (Geng et al., 2025) emphasize robustness, data synthesis, and limited multimodal inputs.
In parallel, reinforcement-learning–based frameworks push the boundaries of search adaptivity and depth. Search-R1 (Jin et al., 2025), R1-Searcher (Song et al., 2025), ReSearch (Chen et al., 2025), Pangu DeepDiver (Shi et al., 2025), and DeepResearcher (Zheng et al., 2025a) train persistence and decomposition skills through reward shaping. ASearcher (Gao et al., 2025) extends this direction with large-scale asynchronous reinforcement learning, decoupling exploration from evaluation to unlock long-horizon search over dozens of hops. Complementary agentic systems such as Open Deep Search (Alzubi et al., 2025), AutoAgent (Tang et al., 2025), and Search-o1 (Li et al., 2025c) further democratize and automate search-enhanced reasoning.
Multimodal Web Search and MM-BrowseComp.
Beyond text-only browsing, MMSearch (Jiang et al., 2024) and MMSearch-R1 (Wu et al., 2025) introduce image-grounded queries, but many tasks still reduce multimodality to entity matching rather than genuine visual reasoning. MM-BrowseComp (Li et al., 2025b) addresses this gap with a 224-question benchmark designed for multimodal browsing agents where key evidence may be embedded in images or videos encountered during search. The benchmark provides a verified checklist for fine-grained analysis of multimodal dependencies and reasoning paths, and reveals that even top models (e.g., OpenAI o3 with tools) achieve only $\sim\!29\%$ accuracy, underscoring substantial headroom for native multimodal reasoning within browsing pipelines. Our setting is closely related to, and concurrent with, MM-BrowseComp, but differs in focusing on sparse visual cues that must be transformed into targeted queries rather than predominantly entity-level matching.
Thinking with Images.
A growing body of work argues that multimodality should go beyond tool orchestration to think with images. Methods such as V* (Wu & Xie, 2024) and instruction-guided masking (Zheng et al., 2024) improve visual grounding inside LLMs. RL-style training further enforces image-first reasoning: GRIT (Fan et al., 2025), DeepEyes (Zheng et al., 2025b), Pixel Reasoner (Su et al., 2025a), and Visionary-R1 (Xia et al., 2025) encourage explicit visual deliberation before action. Visual Agentic Reinforcement Fine-Tuning (Liu et al., 2025), OpenThinkIMG (Su et al., 2025b), and One-RL-to-See-Them-All (Ma et al., 2025) propose unified training frameworks for multimodal agents, while broader systems like PyVision (Zhao et al., 2025) and Thyme (Zhang et al., 2025) generalize the paradigm.
Positioning and Summary.
Taken together, web-browsing benchmarks stress persistence and decomposition in open-world settings (Wei et al., 2025; Li et al., 2025a; Tao et al., 2025; Geng et al., 2025), while multimodal benchmarks emphasize perception-heavy reasoning (Johnson et al., 2017; Krishna et al., 2016; Zellers et al., 2019; Mathew et al., 2021). Recent multimodal browsing datasets—including MM-BrowseComp (Li et al., 2025b)—begin to unite these threads but still leave open how sparse visual evidence should drive query formulation and multi-step planning. We formulate the Spatial-Temporal Extrapolation methodology to concretely design challenging VQAs for testing multimodal search abilities. We also stress the importance of grounding and cropping images to allow fine-grained multimodal reasoning.
7 Conclusion
This work introduces MMSearch-Plus, a benchmark that restores the genuinely multimodal pressures of web browsing by requiring agents to extract weak, localized visual cues and propagate them through noisy retrieval to out-of-image facts. Our Spatial–Temporal Extrapolation curation procedure preserves the persistence and difficulty profile of text-only browsing while compelling fine-grained visual reasoning and provenance verification. Evaluating a range of open and closed MLLMs within a model-agnostic search agent framework reveals a substantial headroom: the strongest agent we tested (o3) reached only 15.1% without search and 36.0% with rollout, while a strong open-source model (Qwen-2.5-VL 72B) achieved 0.0% without search and 6.9% with rollout. These results indicate that current systems struggle to (i) reliably read and localize micro-signals (badges, micro-text, uniforms, layout geometry), (ii) decide when and how to crop and reuse visual evidence, and (iii) maintain verifiable chains of evidence across long-horizon tool use. We hope MMSearch-Plus serves as a rigorous stress test and a common yardstick for progress on multimodal browsing.
Limitations.
While diverse, our data inherits biases from public web sources and the domains we sampled; some tasks still under-represent non-English or low-visual-density pages. Our evaluation focuses on general-purpose agents with standard tool access; specialized systems with richer UI control or domain knowledge may behave differently. Finally, we primarily target image–text browsing; videos and dynamic interfaces remain underexplored.
Appendix A Comparison with Previous Benchmarks
Table 2: Comparison of multimodal information–seeking datasets and benchmarks, ordered by release date.
Benchmark $M$ Search Multi‑hop (>>2) Thinking w/ Image Multi‑Image
Text Image Multi‑turn
InfoSeek
Frames
MMSearch
FactualVQA
ChineseSimpleVQA
BrowseComp
BrowseComp‑ZH
Ours
Appendix B Detailed Dataset Statistics
The distributions of primary categories and secondary categories across MMSearch-Plus are reported in Table 3 and 4, respectively.
Table 3: Primary category distribution.
Category Count %
Geography 64 20.6%
Sports 54 17.4%
Academic Research 50 16.1%
Film & TV 40 12.9%
Technology 36 11.6%
Video Games 31 10.0%
Vlog 19 6.1%
Music 17 5.5%
Table 4: Secondary categories.
Secondary Category Count %
technology 29 9.3%
Team Sports and Match Results 28 9.0%
Countries and Cities 20 6.4%
Television 18 5.8%
Paper and Publication Details 17 5.5%
Experimental Data and Findings 16 5.1%
Movies 16 5.1%
Addresses and Locales 15 4.8%
Transportation Infrastructure 13 4.2%
vlog 12 3.9%
Landmarks and Natural Features 10 3.2%
General Sports Events and Competitions 10 3.2%
Methodological Research 9 2.9%
Game Worlds and Characters 8 2.6%
Game Mechanics 8 2.6%
Esports and Competitive Gaming 8 2.6%
Live Performances 6 1.9%
Cultural and Historical Geography 6 1.9%
Individual Sports and Competitions 6 1.9%
Artists and Bands 5 1.6%
Short Films & Documentaries 4 1.3%
Technological Application 4 1.3%
Racket Sports 4 1.3%
Theoretical and Conceptual Studies 4 1.3%
Gaming Culture and Community 4 1.3%
Channel Characteristics 3 1.0%
Songs and Albums 3 1.0%
Esports and Non-traditional Sports 3 1.0%
Video Game Identification and Content 3 1.0%
Motor and Cycling Sports 3 1.0%
Music Production and Equipment 2 0.6%
Product Review & Launch 2 0.6%
Computing & AI 2 0.6%
Space Technology 1 0.3%
Musical Instruments 1 0.3%
Gadgets & Robotics 1 0.3%
Conventions & Events 1 0.3%
Personal Experience 1 0.3%
Film & TV Personalities 1 0.3%
Innovative Consumer Electronics 1 0.3%
Simulation & Modeling 1 0.3%
Wearable Technology 1 0.3%
Lifestyle & Tips 1 0.3%
Total 311 100.0%
Appendix C Results
The accuracies for o3 and Gemini-2.5-Pro across search modes are reported in Table 5.
Table 5: Results on MMSearch+ (accuracy %). Breakdown by search mode.
Model None One Image Full Rollout
Human
o3 (2025-04-16) 15.1 19.3 36.0
Gemini-2.5-Pro 10.6 16.4 23.8
Appendix D Rollout Trajectory Statistics
The statistics for the rollout trajectories of o3 under full rollout search mode are reported in Tables 6,7,891011 and 12.
Table 6: Quantiles per Conversation (Key Metrics)
Metric 25th Pct 50th Pct 75th Pct 90th Pct 95th Pct 99th Pct
Search calls total 2 3 7 16 20 20
Text queries total 5 11 24.5 49 63 83.8
Assistant words 14 21 36 56 67 119.1
Table 7: Quantiles per Conversation — Correct
Metric 25th Pct 50th Pct 75th Pct 90th Pct 95th Pct 99th Pct
Search calls total 1.0 2.0 5.0 8.9 12.0 15.9
Text search calls 1.0 2.0 5.0 8.9 11.5 15.8
Image search calls 0.0 0.0 0.0 1.0 1.0 1.0
Text queries total 5.0 9.5 19.0 31.0 41.9 61.7
Assistant characters 104.8 139.5 186.5 280.9 413.8 499.5
Assistant words 14.8 20.0 29.3 46.9 66.0 79.3
Tool result items 52.0 89.0 185.0 294.8 424.4 566.3
Tool result characters 18197.5 33739.5 79523.5 109980.6 158017.3 226533.7
Tool result words 2316.5 4543.5 9799.5 14974.8 21038.1 32243.8
Table 8: Quantiles per Conversation — Incorrect
Metric 25th Pct 50th Pct 75th Pct 90th Pct 95th Pct 99th Pct
Search calls total 2.0 3.0 9.0 20.0 20.0 20.0
Text search calls 2.0 3.0 9.0 19.0 20.0 20.0
Image search calls 0.0 0.0 1.0 1.0 1.0 2.0
Text queries total 6.0 13.0 31.0 53.6 67.4 87.0
Assistant chars 94.5 147.0 231.5 368.6 439.1 938.2
Assistant words 14.0 22.0 37.0 59.0 69.3 162.4
Tool result items 68.0 136.0 274.0 494.4 649.8 754.6
Tool result chars 27562.5 48914.0 88466.0 190252.8 256395.3 325728.4
Tool result words 3122.0 6263.0 11937.0 25224.8 34574.4 40940.8
Table 9: Means per Conversation by Correctness (selected metrics)
Metric All Correct Incorrect
Total search calls 5.4 3.7 6.4
Text search calls 5.2 3.5 6.1
Image search calls 0.2 0.1 0.3
Text queries total 19.3 14.2 22.2
Assistant words 27.5 25.1 28.9
Tool result items 179.7 135.1 204.9
Table 10: Top Tool Calls by Name
Tool Count
Text search 1613
Image search 75
Table 11: Top Tool Calls — Correct
Tool Count
Text search 397
Image search 16
Table 12: Top Tool Calls — Incorrect
Tool Count
Text search 1216
Image search 59
Appendix E Detailed Error Types Statistics
In Table 13, the manual annotations for a pilot study on the rollout trajectories of o3 on a previous version of our dataset of about 300 data points are provided.
Error Type Percentage
No relevant information found 51.1%
Hallucination 11.5%
Key information not extracted 8.4%
Relevance not verified 6.9%
Failed to understand the image 6.1%
Refused to answer 4.6%
Multiple answers 3.1%
Question misunderstood 3.1%
Search not invoked 2.3%
May require video understanding 1.5%
Summarization error 1.5%
Total 100.0%
Table 13: Error type distribution (minimal merge).