### v2

# MMSearch-Plus: Benchmarking Provenance-Aware Search for Multimodal Browsing Agents

Abstract
Existing multimodal browsing benchmarks often fail to require genuine multimodal reasoning, as many tasks can be solved with text-only heuristics without vision-in-the-loop verification. We introduce MMSearch-Plus, a 311-task benchmark that enforces multimodal understanding by requiring extraction and propagation of fine-grained visual cues through iterative image–text retrieval and cross-validation under retrieval noise. Our curation procedure seeds questions whose answers require extrapolating from spatial cues and temporal traces to out-of-image facts such as events, dates, and venues. Beyond the dataset, we provide a model-agnostic agent framework with standard browsing tools and a set-of-mark (SoM) module, which lets the agent place marks, crop subregions, and launch targeted image/text searches. SoM enables provenance-aware zoom-and-retrieve and improves robustness in multi-step reasoning. We evaluated closed- and open-source MLLMs in this framework. The strongest system achieves an end-to-end accuracy of 36.0%, and integrating SoM produces consistent gains in multiple settings, with improvements up to +3.9 points. From failure analysis, we observe recurring errors in locating relevant webpages and distinguishing between visually similar events. These results underscore the challenges of real-world multimodal search and establish MMSearch-Plus as a rigorous benchmark for advancing agentic MLLMs.
Project page: https://mmsearch-plus.github.io/
footnotetext: *Equal contribution. Corresponding authors: liu.rui2@huawei.com, lpk@cs.hku.hk.
1 Introduction
Figure 1: Three multimodal reasoning paradigms: 1. Without search: an MLLM reasons with its internal knowledge to answer a factual visual question (VQA); 2. Whole image search only: an MLLM combines its internal knowledge and the provided search results of an VQA’s image; 3. MMSearch-Plus agentic framework: an MLLM can call a set of visual and search tools freely to extracts fine-grained visual clues and search with precision. A real example from MMSearch-Plus is provided in Figure 8.
Recent advances in large multimodal language models (MLLMs) have enabled them to act as capable agents that integrate language, vision, and tool use to accomplish complex tasks. A growing body of work studies how MLLMs leverage external search engines—both text- and image-based—to answer information-seeking questions.
Multimodal browsing benchmarks such as MMSearch (Jiang et al., 2024) primarily test a model’s ability to coordinate text and images. However, many MMSearch-style questions can be solved by relatively fixed workflows that require little genuine multimodal reasoning. In particular, MMSearch heavily relies on external image search: the MLLM often plays the role of an orchestrator rather than a visual reasoner. When the search engine retrieves highly relevant images, even a unimodal LLM—without visual perception—can frequently answer by reasoning over the accompanying page text alone. This occurs because a single strong image search can surface pages whose surrounding text already contains the answer. In such cases, the image search tool and the MLLM are partially interchangeable as sources of image-related information; the only distinctly multimodal step is often image–source cross-validation, a narrow slice of multimodal reasoning.
Empirically, MMSearch requires short reasoning chains and achieves relatively high model scores: GPT-4o exceeds 60% on MMSearch (end-to-end VQA), and MMSearch-R1-7B reports 54.6% average accuracy (Wu et al., 2025). By contrast, recent text-only browsing benchmarks emphasize persistence and creative, multi-step search. BrowseComp (Wei et al., 2025) targets web navigation for hard-to-find, entangled information; on this task, GPT-4o and GPT-4.5 each score below 1% in direct-answer settings and under 2% even with browsing tools (Wei et al., 2025). “Deep Research” reports 51.5% on BrowseComp, though it was trained for the benchmark and public implementation details remain limited, while open-source agents such as WebSailor reach roughly 12% after substantial engineering (Li et al., 2025a). Taken together, these results suggest that current multimodal benchmarks may not yet match the difficulty of complex text-only browsing, motivating a BrowseComp-style multimodal benchmark with similarly demanding reasoning.
What makes a challenging benchmark for multimodal browsing agents.
Building on these insights, we argue that a strong multimodal benchmark should: (i) require fine-grained, exhaustive visual reasoning that compels models to mine subtle, localized cues rather than rely on a single salient (“dominant”) entity that occupies most of the image; and (ii) enforce source verification under retrieval noise—i.e., discriminating authentic sources when image results are noisy or conflicting and validating images embedded in webpages. Concretely, the benchmark should stress robust cross-modal evidence gathering, careful verification, and resilience to near-duplicates. For example, two football matches from adjacent seasons may share nearly identical visuals (kits, stadium angles). Correctly answering requires inspecting localized markers (e.g., patch details, sponsor variants, scoreboard typography) and cross-checking dates and venues against trustworthy sources. In real applications, difficult multimodal information-seeking tasks exhibit three recurring challenges that our benchmark targets:
  1. 1.
    Noisy or conflicting retrieval. Image search may return non-relevant or contradictory results; agents must identify authentic sources via fine-grained multimodal understanding and cross-source consistency checks.
  2. 2.
    Exhaustive, part-based visual reasoning. When holistic matching fails, agents must reason over parts and subregions—often iteratively cropping and re-searching—to piece together fragmented clues that identify the target entity or event.
  3. 3.
    Professional multimodal handling and long, tool-augmented chains. Agents must extract answers from heterogeneous media (images, documents, videos) across the web, sustain long reasoning/tool-use chains, and invoke visual tools programmatically (e.g., cropping, OCR) when helpful, akin to advanced agents such as o3.
Our work, MMSearch-Plus, aims to bridge the gap by combining the persistence and high-reasoning demands of BrowseComp with truly multimodal workflows. Unlike (Geng et al., 2025), where images often collapse to a single identifiable entity, images in our benchmark contain multiple subtle visual signals that must be connected throughout the entire reasoning process. To systematically curate hard multimodal search questions, we introduce Spatial-Temporal Extrapolation, a principled curation procedure that seeds questions whose answers require extrapolating from spatial (localized visual cues, layouts, micro-text, uniforms, signage) and temporal signals (broadcast overlays, seasonal context, historical rosters/schedules) to identify events, dates, or locations not explicitly present in the prompt or image. This forces models to extract fine-grained cues, propagate them through iterative (image/text) searches, and verify provenance amidst retrieval noise—directly addressing the challenges above.
MMSearch-Plus targets realistic scenarios where answers cannot be read off from the prompt or the image alone. For example, given a screenshot from an esports broadcast, the agent may need to infer the match date—absent from both the query and the image—by integrating subtle overlays with external evidence. Similar patterns arise in sports, geolocation, and lifestyle vlogs, domains we explicitly include to move beyond typical web imagery.
Our key contributions can be summarized as follows:
  • We propose MMSearch-Plus, a challenging benchmark that integrates fine-grained multimodal reasoning (with sparse, localized visual signals) and external tool use, while maintaining the difficulty profile of text-based browsing tasks like BrowseComp (Wei et al., 2025).
  • We introduce the curation procedure Spatial-Temporal Extrapolation ($\xi$ 2.1) to construct difficult multimodal search questions that require extrapolating from localized spatial cues and temporal signals to out-of-image facts.
  • We develop a general, model-agnostic search agent framework that interleaves text/image search with a Set-of-Mark zoom-in pathway ($\xi$ 2.5). SoM supplies human-verified marks and region-seeded retrieval (zoom_in/image_search), enabling sustained fine-grained reasoning; we ablate SoM to quantify how cropping and thinking with images affect current MLLMs.
A concurrent work, BrowseComp-VL (Geng et al., 2025), increases difficulty mainly by expanding the text search space in a BrowseComp-like manner, while the image often reduces to a single identifiable entity that is straightforwardly searchable and quickly disposable. Another concurrent work, MM-BrowseComp (Li et al., 2025b), is related. Our contributions differ in (a) data sources and (b) a general search framework that also enables thinking with images via Set-of-Mark.
2 Method
2.1 Spatial-Temporal Extrapolation
Summary. We test an agent’s ability to infer facts that are not literally present in the pixels or the instant shown, requiring it to recover the broader event by combining partial visual cues with retrieval and reasoning across time and space. This emphasis discourages purely perceptual shortcutting and directly measures uncertainty-aware planning, precise event localization, and evidence-backed browsing.
A central challenge in BrowseComp-like benchmarks is the ballooning intermediate search space induced by soft, fuzzy constraints, which demands non-trivial cross-validation to identify the correct target. Rather than remixing text-only corpora, we construct multimodal problems that naturally expand this search space: from a handful of visual fragments tied to real-world events (in the spirit of GeoGuessr AB (2013)), the agent must hypothesize the underlying source event and verify it against retrieved evidence. Difficulty is modulated by the richness of visual and textual cues; even a single crop can widen the candidate set dramatically, eliciting iterative hypothesis–test–refine behavior and long trajectories that require robust context tracking.
Figure 2: Examples of spatial and temporal relations for question curation.
Once the source event is pinned down, our questions probe metadata and support multi-hop reasoning. To elevate difficulty, we introduce Spatial-Temporal Extrapolation: instead of querying what is directly visible, we ask what is contextually implied but physically absent, compelling reasoning beyond the frame and moment to reconstruct the broader situation. Spatial extrapolation targets unseen entities (e.g., off-screen participants, individuals facing away, occluded signage), while temporal extrapolation probes events immediately before or after the depicted moment (e.g., lineups, the next play/goal, or the subsequent segment/episode). Solving these requires agents to (i) precisely localize the event (time, match, or episode), and (ii) retrieve and integrate wider contextual knowledge from diverse sources to produce an evidence-grounded answer.
2.2 Data Collection
Building on the Spatial–Temporal Extrapolation framework, we instantiated the data collection process as follows.
We began with a set of predefined categories and generated search keywords using a mix of LLM assistance and manual curation. Data were sourced from publicly accessible video-sharing platforms and open-access scholarly repositories (i.e., arXiv). Queries for consumer video content (e.g., Vlog) were issued to video platforms, while those in the Acad. category were issued to arXiv to retrieve relevant research papers (PDFs). All materials are used only for research. We also mask sensitive content (e.g., faces and vehicle license plates). In our pipeline, both videos and papers are treated as events.
For videos, human annotators manually selected keyframes suitable for constructing questions. The keyframes underwent anonymization for sensitive content. For papers, annotators extracted figures or tables, capturing them via screenshots. A screenshot was considered suitable if it (i) contained vague or noisy information that challenged perception, (ii) depicted less familiar entities requiring external retrieval to answer associated questions, and (iii) could not be searched with Google image search directly. This ensures that most, if not all, screenshots are not in the training dataset, and the essential information for answering the question is neither in the training corpora nor directly retrievable via a direct image search.
Filtering.
Although the collection focused on new or rarely seen events, we observed that certain closed-source models (e.g., GPT-4o, GPT-5, Gemini-2.5-Pro) could occasionally solve questions without external search. To counteract this, we employed an adversarial filtering procedure to increase difficulty. Specifically, annotators removed or obfuscated questions answerable without retrieval through the following methods:
  • Cross-validation: Annotators designed problems outside their own knowledge base, then verified solvability by testing on at least two closed-source MLLMs.
  • Image perturbation: Critical visual subregions were blurred or masked.
  • Iterative refinement: Questions that remained trivially solvable were discarded.
2.3 Dataset Statistics
Statistic Number
Total questions w/ images 311
Primary categories 8
Secondary categories 43
Unique images 441
Unique questions 279
Unique answers 306
Maximum question length 42
Maximum answer length 20
Average question length 12.5
Average answer length 3.7
Difficulty distribution
- Easy 94 (30.2%)
- Difficult 217 (69.8%)
Figure 3: (a) Distribution of primary categories in MMSearch-Plus, with light and dark inner sectors representing the easy and difficult splits within each category. (b) Summary of statistics.
Figure 3 depicts the distribution of MMSearch-Plus tasks across eight primary categories, which is approximately balanced. Each task includes at least one image, with additional images provided when a single view is insufficient. The benchmark primarily comprises short-answer questions, with answers averaging 3.7 words. In hindsight, we labeled the difficulty level of a task as Easy if it can be answered correctly without search or with only image search by o3 or Gemini-2.5-Pro. Otherwise, the task is labeled as Difficult (interchangeably Hard).
2.4 Benchmark Metric and Evaluation
We adopt accuracy as the primary evaluation metric. For each question, a set of acceptable answers is provided. We use LLM-as-a-judge to compare model predictions against this set, deeming an answer correct if it matches any acceptable variant. Validation through manual inspection showed complete agreement between human judgment and the assessment of GPT-4o.
2.5 Search Agent Framework
Overview.
We implement a model-agnostic web agent that interleaves text search, image search, and a zoom-in tool. Ranked results are obtained via SerpApi ; the surrounding framework (i.e., the system that includes the MLLM agent) performs lightweight post-processing to summarize pages, relate images to the query, and manage multi-round context. To reduce latency and token cost across repeated steps/evaluations, the framework caches image-search returns (ranked URLs/thumbnails) and their MLLM summaries per benchmark image. The framework maintains a threaded conversation state (tool calls, cropped views, summaries, and hypotheses) to support long-horizon, contextualized reasoning. It is compatible with sub-image cropping and sub-image search for fine-grained analysis.
Refer to caption Q: Who assisted this goal? A: Philipp Lahm
Figure 4: Example of an SoM-overlaid task image.
Text and image search.
At any step, the agent may invoke text or image search. SerpAPI returns ranked candidates, from which the framework forwards a small set of top results to the model (images are summarized and cached; webpages are summarized). Retrieved webpages are summarized by an MLLM (Gemini in our runs) into two fields:
  • web_info: task-conditioned semantic summaries of page content that may support answering the current question;
  • related_info: evidence linking a result’s thumbnail/lead image(s) to the query image (e.g., matching signage, layouts, or micro-text).
These summaries are designed to compress interaction history for use within limited MLLM context windows; metadata such as URL/title is readily handled by SerpAPI. All retrieved and generated artifacts are integrated into the conversation state to enable contextualized evaluation across turns.
Zoom-in via Set-of-Mark (SoM).
To enable precise, provenance-aware inspection, we augment the agent with a Set-of-Mark (SoM) module (Yang et al., 2023). In this work, we refer to SoM as a list of human-verified bounding boxes per task image (examples in Fig. 9). In the first turn, we present both the raw image and an overlaid version with box outlines and indices, avoiding occlusion while allowing cross-reference.
Table 1: End-to-end results on the MMSearch+ benchmark, across search modes. All numbers are accuracy (%). Darker blue means higher accuracy. The best and second-best methods under each setting are shown in bold and underline, respectively.
Model / Search Mode Avg By Category Difficulty
Geo. Sports Acad. Film/TV Tech Games Vlog Music Easy Hard
Human
Browser 22.8 20.3 25.9 20.0 25.0 19.4 16.1 31.6 35.3 34.0 18.0
Closed-source
o3 (2025-04-16)
Without Search 15.1 31.2 14.8 6.0 17.5 13.9 3.2 5.3 11.8 50.0 0.0
Image Search 19.3 28.1 14.8 18.0 30.0 22.2 3.2 5.3 17.6 63.8 0.0
Full Rollout 36.0 35.9 24.1 50.0 42.5 44.4 16.1 42.1 29.4 54.3 28.1
Full Rollout + SoM 37.7 45.3 29.6 46.0 45.0 45.7 16.1 26.3 29.4 62.8 26.9
GPT-5
Without Search 10.3 21.9 7.4 4.0 7.5 8.3 0.0 5.3 15.8 27.7 2.8
Image Search 16.4 25.0 11.1 14.0 22.5 19.4 3.2 0.0 29.4 50.0 1.8
Gemini-2.5-Pro
Without Search 10.6 15.6 11.1 6.0 12.5 13.9 0.0 15.8 5.9 35.1 0.0
Image Search 16.4 26.6 11.1 18.0 20.0 16.7 3.2 0.0 23.5 54.3 0.0
Full Rollout 23.8 39.1 14.8 12.0 27.5 33.3 6.5 26.3 29.4 46.8 13.8
Full Rollout + SoM 27.7 40.6 22.2 24.0 25.0 33.3 19.4 15.8 29.4 54.3 16.1
Open-source
Qwen-2.5-VL-72B
Without Search 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Image Search 13.5 20.3 7.4 18.0 17.5 11.1 3.2 0.0 23.5 41.5 1.4
Full Rollout 6.1 9.5 7.4 4.0 5.0 2.8 3.2 5.3 11.8 17.0 1.4
Full Rollout + SoM 7.1 10.9 3.7 4.0 10.0 5.6 6.5 5.3 11.8 18.1 2.3
3 Experiment
Evaluated Models.
We evaluate both closed-source and open-source MLLMs on our benchmark. Closed-source models include Gemini-2.5-Pro (hereafter Gemini(Team & DeepMind, 2025), o3 (OpenAI, 2025b) and GPT-5 (OpenAI, 2025a), representing the current frontier of proprietary multimodal systems. On the open-source side, we focus on Qwen-2.5-VL-72B-Instruct (hereafter Qwen(Bai et al., 2025). Expert human annotators were instructed to work with a Chrome browser for at least 10 minutes for each task, and capped at 20 minutes.
Implementation Details.
To ensure a fair comparison, we standardize the search-and-reasoning pipeline across all models. We consider three search modes:
  1. (a)
    Without Search. The MLLM answers using only the question and image inputs.
  2. (b)
    Image Search. For each benchmark image, the top-10 image-search results are summarized by Gemini and cached. The MLLM then receives the question, the images, and these summaries as input. For multi-image questions, we append each image’s summaries to the model context.
  3. (c)
    Full Rollout. The model runs our search-agent framework and may call search tools as needed. Image-search settings match those in Image Search. For text search, we retrieve the top-5 websites by default. The process is capped at 20 search rounds (combining text and image search) to balance efficiency and coverage. In each text-search round, the model generates 3–5 refined queries; in each image-search round, it issues exactly one query. For Qwen, pilot studies showed that more than 10 rounds yield no further gains, so we cap its search at 10 rounds.
  4. (d)
    Full Rollout + SoM. Images overlayed with Set-of-Mark are provided to MLLMs. In addition to capabilities provided in Full Rollout, models can (a) call the zoom-in tool to examine a subregion of an image, (b) call the image search tool to search with the subregion. The maximum number of tool calls is 20.
4 Results
The results of MLLMs across search modes are presented in Table 1.
Under No Search, open-source Qwen scored 0.0%, confirming that MMSearch+ resisted parametric shortcuts. Closed-source models remained far from saturation (GPT-5 10.3, Gemini-2.5-Pro 10.6, o3 15.1). With Image Search (o3 +4.2, GPT-5 +6.1, Gemini +5.8; Qwen +13.5), models gained from coarse disambiguation, yet multi-hop composition remained the main blocker. With iterative image and text search, Full Rollout produced large gains: o3 36.0 (+20.9 over no-search) and Gemini 23.8 (+13.2), while Qwen regressed to 6.1 (-7.4 over image-search). This indicates that the external information retrieved from search tools is essential to answering the questions correctly. Notably, easy split accuracy did not improve from Image Search to Full Rollout. We attribute this to two factors:
  1. 1.
    Label vs. retrieval difficulty. In MMSearch+, “easy” meant higher likelihood of being in-model or solvable by a direct image hop, not that multi-step retrieval was trivial.
  2. 2.
    Policy overhead and distractors. Full rollout added branching (tool choice, query rewriting). For easy items solvable by a single image search, extra hops often injected distractors (off-topic pages, SEO noise), increasing drift from the correct evidence.
Overall, rollout helped where progressive narrowing was necessary (academic, tech, media queries), but on easy items extra freedom reduced precision.
Adding SoM (Full Rollout + SoM) gave consistent, targeted gains: o3 37.7 (+1.7), Gemini 27.7 (+3.9), Qwen +1.0 over Full Rollout. Set-of-Mark anchored reasoning on fine-grained visual cues (micro-text, logos, key entities) before expanding the search space.
5 Analysis
We present a list of observations (Section 5.1) and an error analysis (Section 5.2).
5.1 Observations
Figure 5: Stepwise tool distribution for Gemini, o3 and Qwen under Full Rollout+SoM, capped at Step $12$ for visualization. We use search round and step interchangeably, and an MLLM can make multiple search queries in one round (details in $\xi$ 3).
Closed-source models retained broad geographic knowledge.
They often recognized rare locales and supplied plausible context; geography tasks were comparatively strong. Without search tools, o3 achieved 31.2%, while human annotators with a standard browser only achieved 20.3%. On other categories like sports and Vlog, the human baseline usually outperformed the tested MLLMs. We name this phenomenon the different comparative advantages of humans and MLLMs.
o3 exhibited strong long-context coordination, while Gemini leveraged fewer search calls.
On the one hand, o3 sustained 10+10{+} search rounds and reasoned over $\sim 50{+}$ retrieved items without excessive redundancy, contributing to its rollout gains. On the other hand, Figure 5 (left) shows that Gemini always answered on or before Step 9. However, more tool calls did not correlate with trajectory correctness: Figure 14 shows that incorrect trajectories typically contained more search calls than correct ones.
Format correctness of closed-source models’ tool calls, and Qwen’s tool-use instability.
All the closed-source models consistently produced valid tool calls. For Qwen, we observed $421$ invalid image-search calls across $45$ of $311$ tasks, inflating retries without quality improvements.
More text search calls than image search calls, and zoom-in operations were not always followed by a sub-image search call.
In Figure 5, all models made more text-search calls than image-search calls. In addition, MLLMs sometimes zoomed in on the SoM-cropped image subregion only to see it clearly, without making a subsequent image search using the sub-image as the query.
5.2 Error Analysis
Figure 6: Distribution of error types. Annotated on Gemini-2.5-Pro (full rollout).
We categorize failure cases using a manually curated taxonomy and report the distribution over categories in Figure 6. Labels are assigned by the authors (single dominant label per erroneous prediction) for Gemini-2.5-Pro under the full-rollout setting. Examples of incorrect task trajectories are provided in Figure 7.
High-frequency categories.
The two most prevalent failure modes are no relevant information found and hallucination, responsible for $51.1\%$ and $11.5\%$ of errors, respectively ($62.6\%$ combined). The former typically reflects a retrieval failure: noisy or overly specific queries miss on-topic pages. The latter reflects vision-driven misattribution: the model latches onto an incorrect event from internal knowledge and either does not invoke search or discounts contradictory evidence in the retrieved pages.
Other observed categories.
We also observe failures from misinterpreting the multimodal context; hallucinations grounded in either images or parametric knowledge; refusals or multi-answer outputs when evidence is insufficient to disambiguate a single event; failing to invoke search (falling back on outdated priors); and errors when summarizing relevant webpages.
Qualitative takeaways and implications.
Two recurring patterns emerge: (i) there is no single “golden path”—agents can succeed via varied orders of search/crop/verify; and (ii) stronger multimodal internal knowledge narrows the practical search space, enabling targeted verification over broad exploration. While augmenting multimodal internal knowledge can reduce exploration overhead, end-to-end reliability still depends on robust verification and tighter cross-modal grounding.
Figure 7: Representative error case studies. Each row illustrates an example from a high-frequency error type (Figure 6). Panels show the prompt, GT answer, and a shortened rollout. From top to bottom: (a) Key info not extracted. (b) Relevance not verified. (c) Question misunderstood.
6 Related Work
Classical vision–language benchmarks probe compositional and commonsense reasoning in images and text (Johnson et al., 2017; Krishna et al., 2016; Zellers et al., 2019; Mathew et al., 2021), with recent suites extending to long videos and interleaved modalities (Wu et al., 2024; Li et al., 2024; Hu et al., 2025; Jeong et al., 2025). Concurrent work on web agents studies planning, browsing, and evidence gathering in open worlds (Krishna et al., 2024; Wei et al., 2025; Geng et al., 2025; Li et al., 2025a; Tao et al., 2025). Multimodal web search benchmarks begin to couple vision with browsing, but often emphasize entity matching over fine-grained visual reasoning (Jiang et al., 2024; Wu et al., 2025). Closest to our setting, MM-BrowseComp targets image/video evidence encountered during search and reveals substantial headroom even for frontier models (Li et al., 2025b). A complementary line promotes “thinking with images” via improved grounding and RL-style training to elicit image-first reasoning (Wu & Xie, 2024; Zheng et al., 2024; Fan et al., 2025; Zheng et al., 2025b; Su et al., 2025a; Xia et al., 2025). We differ by targeting sparse, spatial–temporal visual cues that must be transformed into targeted queries and multi-step plans, providing a focused stress test for multimodal browsing agents.
7 Conclusion
This work introduces MMSearch-Plus, a benchmark that restores the genuinely multimodal pressures of web browsing by requiring agents to extract weak, localized visual cues and propagate them through noisy retrieval to out-of-image facts. Our Spatial–Temporal Extrapolation curation procedure preserves the persistence and difficulty profile of text-only browsing while compelling fine-grained visual reasoning and provenance verification. Evaluating a range of open and closed MLLMs within a model-agnostic search agent framework reveals a substantial headroom: the strongest agent we tested (o3) reached only 15.1% without search and 36.0% with rollout, while a strong open-source model (Qwen-2.5-VL-72B-Instruct) achieved 0.0% without search and 6.9% with rollout. These results indicate that current systems struggle to (i) reliably read and localize micro-signals (badges, micro-text, uniforms, layout geometry), (ii) decide when and how to crop and reuse visual evidence, and (iii) maintain verifiable chains of evidence across long-horizon tool use. We hope MMSearch-Plus serves as a rigorous stress test and a common yardstick for progress on multimodal browsing.
Appendix A Use of LLMs
The manuscript text was polished for fluency using Grammarly after the authors completed a full draft. The authors then manually reviewed all suggestions to ensure their appropriateness. In addition, ChatGPT was used to generate the task illustration in Figure 1 (upper left), based on the real image in Figure 8.
Appendix B Example from MMSearch-Plus
Figure 8: Example MMSearch-Plus item. Given a 2025 concert photo and the query “What was the singer’s performance time?”, the agent’s golden path extracts localized cues—micro-text/lyrics, a crop of the singer, festival/brand signage, and a distinctive stage prop—and issues targeted image/text searches that (i) identify the artist/outfit, (ii) resolve the event (e.g., a specific festival and city/date), and (iii) retrieve and cross-validate the official schedule to obtain the exact performance time. The answer is not explicitly written in the prompt or the image, underscoring MMSearch-Plus’s emphasis on fine-grained multimodal reasoning with provenance checks under retrieval noise.
Appendix C Comparison with Previous Benchmarks
Table 2: Comparison of multimodal information–seeking datasets and benchmarks, ordered by release date.
Benchmark/ Dataset $M$ Search
Multi-hop
(>>2)
Thinking
w/ Image
Multi-
Image
Text Image
Multi-
turn
InfoSeek (Chen, 2023)
FRAMES (Krishna et al., 2024)
MMSearch (Jiang et al., 2024)
FactualVQA(Wu et al., 2025)
ChineseSimpleVQA (Gu et al., 2025)
BrowseComp (Wei et al., 2025)
BrowseComp-ZH (Zhou et al., 2025)
BrowseComp-VL (Geng et al., 2025)
MM-BrowseComp (Li et al., 2025b)
Ours
Appendix D Related Work
We present a more comprehensive literature review in the Appendix.
Multimodal Reasoning Benchmarks.
Classical vision-and-language benchmarks such as CLEVR (Johnson et al., 2017), VCR (Zellers et al., 2019), Visual Genome (Krishna et al., 2016), and InfographicVQA (Mathew et al., 2021) established core evaluations for compositional reasoning, commonsense inference, and visual question answering. Recent efforts extend these settings to longer contexts and interleaved modalities, including LongVideoBench (Wu et al., 2024), MVBench (Li et al., 2024), Video-MMMU (Hu et al., 2025), and VideoRAG (Jeong et al., 2025), which highlight long-context understanding and retrieval augmentation. Other work such as MIRB (Zhao et al., 2024), pixel-level VCR (Zhang et al., 2024), and ReasonMap (Feng et al., 2025) aim to connect perception with explicit reasoning. However, these datasets are primarily self-contained and do not require dynamic interaction with the open web or tool-mediated search.
Web Browsing and Search-Enhanced Agents.
A large body of work explores language models as web agents that plan, browse, and gather evidence. Earilier works like FRAMES (Krishna et al., 2024) focus on multi-hop QAs on Wikipedia-like data. More recently, benchmarks such as BrowseComp (Wei et al., 2025) and its Chinese variant BrowseComp-ZH (Zhou et al., 2025) test multi-step persistence in text-heavy web navigation, where even frontier models struggle. Open-source agents including WebSailor (Li et al., 2025a), WebShaper (Tao et al., 2025), and WebWatcher (BrowseComp-VL) (Geng et al., 2025) emphasize robustness, data synthesis, and limited multimodal inputs.
In parallel, reinforcement-learning–based frameworks push the boundaries of search adaptivity and depth. Search-R1 (Jin et al., 2025), R1-Searcher (Song et al., 2025), ReSearch (Chen et al., 2025), Pangu DeepDiver (Shi et al., 2025), and DeepResearcher (Zheng et al., 2025a) train persistence and decomposition skills through reward shaping. ASearcher (Gao et al., 2025) extends this direction with large-scale asynchronous reinforcement learning, decoupling exploration from evaluation to unlock long-horizon search over dozens of hops. Complementary agentic systems such as Open Deep Search (Alzubi et al., 2025), AutoAgent (Tang et al., 2025), and Search-o1 (Li et al., 2025c) further democratize and automate search-enhanced reasoning.
Multimodal Web Search and MM-BrowseComp.
Beyond text-only browsing, MMSearch (Jiang et al., 2024) and MMSearch-R1 (Wu et al., 2025) introduce image-grounded queries, but many tasks still reduce multimodality to entity matching rather than genuine visual reasoning. MM-BrowseComp (Li et al., 2025b) addresses this gap with a 224-question benchmark designed for multimodal browsing agents where key evidence may be embedded in images or videos encountered during search. The benchmark provides a verified checklist for fine-grained analysis of multimodal dependencies and reasoning paths, and reveals that even top models (e.g., OpenAI o3 with tools) achieve only $\sim\!29\%$ accuracy, underscoring substantial headroom for native multimodal reasoning within browsing pipelines. Our setting is closely related to, and concurrent with, MM-BrowseComp, but differs in focusing on sparse visual cues that must be transformed into targeted queries rather than predominantly entity-level matching.
Thinking with Images.
A growing body of work argues that multimodality should go beyond tool orchestration to think with images. Methods such as V* (Wu & Xie, 2024) and instruction-guided masking (Zheng et al., 2024) improve visual grounding inside LLMs. RL-style training further enforces image-first reasoning: GRIT (Fan et al., 2025), DeepEyes (Zheng et al., 2025b), Pixel Reasoner (Su et al., 2025a), and Visionary-R1 (Xia et al., 2025) encourage explicit visual deliberation before action. Visual Agentic Reinforcement Fine-Tuning (Liu et al., 2025), OpenThinkIMG (Su et al., 2025b), and One-RL-to-See-Them-All (Ma et al., 2025) propose unified training frameworks for multimodal agents, while broader systems like PyVision (Zhao et al., 2025) and Thyme (Zhang et al., 2025) generalize the paradigm.
Positioning and Summary.
Taken together, web-browsing benchmarks stress persistence and decomposition in open-world settings (Wei et al., 2025; Li et al., 2025a; Tao et al., 2025; Geng et al., 2025), while multimodal benchmarks emphasize perception-heavy reasoning (Johnson et al., 2017; Krishna et al., 2016; Zellers et al., 2019; Mathew et al., 2021). Recent multimodal browsing datasets—including MM-BrowseComp (Li et al., 2025b)—begin to unite these threads but still leave open how sparse visual evidence should drive query formulation and multi-step planning. We formulate the Spatial-Temporal Extrapolation methodology to concretely design challenging VQAs for testing multimodal search abilities. We also stress the importance of grounding and cropping images to allow fine-grained multimodal reasoning.
Appendix E Detailed Dataset Statistics
Table 3: Primary category distribution.
Category Count %
Geography 64 20.6%
Sports 54 17.4%
Academic Research 50 16.1%
Film & TV 40 12.9%
Technology 36 11.6%
Video Games 31 10.0%
Vlog 19 6.1%
Music 17 5.5%
The distributions of primary categories across MMSearch-Plus are reported in Table 3.
Appendix F Set-of-Mark Implementation Details
Figure 9 shows a list of task images overlaid with Set-of-Mark bounding box annotations from MMSearch-Plus.
In our context, set-of-mark for a task image consists of bounding boxes of key visual entities. Each box has a unique index. In the first user turn, we present both the raw image and an overlaid version with box outlines and indices, avoiding occlusion while allowing cross-reference. The framework exposes:
  • zoom_in(index): return a cropped sub-image for localized reasoning;
  • image_search(index): launch an image search seeded by the cropped region (compatible with the standard image-search pathway).
We include SoM because baseline MLLMs were unreliable at proposing visually precise crops from scratch. Providing accurate marks reduces the burden on detection/localization and lets models focus on reading, matching, and reasoning. While automatic mark generation (e.g., strong open-vocabulary detectors) is conceivable, robust coverage is challenging due to long-tail entities common in our benchmark; therefore MMSearch-Plus uses human-curated boxes.
Refer to caption Q: What is the title of this paper? A: MUCAR: Benchmarking Multilingual Cross-Modal Ambiguity Resolution for Multimodal Large Language Models
Refer to caption Q: What is the name of the previous level? A: Exercise Hall
Refer to caption Q: What is the title of this video? A: Who Can Identify as a Native American?
Refer to caption Q: In Paris Fashion Week, what brand did the person in the image represent? In which year did the event occur, and what is the event’s name? A: StreetStyle Winter 2025 Hermes
Figure 9: Examples of task images overlaid with Set-of-Mark bounding box annotations from MMSearch-Plus.
Appendix G Evaluation Details
We initially tested both GPT-4o and Gemini-2.5-Pro as judges and found their outputs consistent. The subsequent experiments used GPT-4o. To further support reproducibility, we optionally provide lightweight rule-based verification methods.
Appendix H Agent Behavior Analysis
This appendix presents detailed visualizations of tool usage patterns across the three multimodal agents: Gemini, o3, and Qwen. All figures show episodes capped at 20 steps for readability.
H.1 Episode Raster Plots
Figure 10 shows the tool usage patterns for the 40 longest episodes from each agent. Each row represents one episode, sorted by length (longest at top). Columns represent sequential steps, with colors indicating tool types: text search, image search, zoom-in, and answer. Dark blue strips represent an MLLM did not fully utilize 20 rounds and stopped early.
Figure 10: Episode raster plots showing tool usage patterns for the 40 longest episodes per agent. Each row is an episode (sorted by length), each column is a step.
H.2 Tool Transition Matrices
Figure 11 shows the Markov transition probabilities between tool types.
[…]
H.3 Zoom-In Behavior Analysis
Figures 12 and 13 analyze zoom-in behavior patterns across agents. o3 shows the most aggressive zooming behavior, while Qwen and Gemini are more conservative.
Figure 12: Distribution of first zoom-in step positions. o3 tends to zoom earlier in episodes compared to Gemini and Qwen.
Figure 13: Distribution of zoom-in frequency per episode. o3 exhibits the highest zoom-in frequency, with some episodes containing 10+ zoom operations.
Appendix I Rollout Trajectory Statistics
Figure 14: Reasoning-trajectory statistics. Rows: MLLMs (o3, Gemini, Qwen from top to down). Columns: left—distribution of image search calls and text search calls per trajectory; right—number of words in an MLLM’s response versus the number of search calls, stratified by correctness. Statistics collected in full rollout mode.
Figure 14 shows the distribution of number of search calls for correct and incorrect task trajectories (left) and the relation of the number of words in a model’s trajectory (excluding the words for tool calls) vs. the number of search calls (right) under the Full Rollout setting.
Appendix J Prompts
System Prompt