### v3

# MMSearch-Plus: Benchmarking Provenance-Aware Search for Multimodal Browsing Agents

Abstract
Existing multimodal browsing benchmarks often fail to require genuine multimodal reasoning, as many tasks can be solved with text-only heuristics without vision-in-the-loop verification. We introduce MMSearch-Plus, a 311-task benchmark that enforces multimodal understanding by requiring extraction and propagation of fine-grained visual cues through iterative image–text retrieval and cross-validation under retrieval noise. Our curation procedure seeds questions whose answers require extrapolating from spatial cues and temporal traces to out-of-image facts such as events, dates, and venues. Beyond the dataset, we provide a model-agnostic agent framework with standard browsing tools and a set-of-mark (SoM) module, which lets the agent place marks, crop subregions, and launch targeted image/text searches. SoM enables provenance-aware zoom-and-retrieve and improves robustness in multi-step reasoning. We evaluated closed- and open-source MLLMs in this framework. The strongest system achieves an end-to-end accuracy of 36.0%, and integrating SoM produces consistent gains in multiple settings, with improvements up to +3.9 points. From failure analysis, we observe recurring errors in locating relevant webpages and distinguishing between visually similar events. These results underscore the challenges of real-world multimodal search and establish MMSearch-Plus as a rigorous benchmark for advancing agentic MLLMs.
1 Introduction
Large multimodal language models (MLLMs) increasingly act as agents that integrate vision, language, and web search to answer information-seeking questions. Several recent benchmarks study this capability by pairing images with web browsing or image search tools. However, despite their multimodal interface, existing multimodal browsing benchmarks rarely demand sustained, fine-grained visual reasoning or long-horizon, tool-augmented search.
This paper introduces MMSearch-Plus, a new benchmark that closes this gap. MMSearch-Plus is designed to require (i) localized, exhaustive visual reasoning over subtle cues, (ii) robust verification under noisy or conflicting retrieval, and (iii) multi-step tool use that interleaves text and image search with region-level visual analysis. To support this, we contribute:
A challenging multimodal browsing benchmark that matches the long-horizon difficulty of text-based browsing tasks (e.g., BrowseComp) while adding visual reasoning requirements that cannot be bypassed by a single strong image search.
Spatial–Temporal Extrapolation, a principled curation procedure that constructs hard questions by requiring models to extrapolate from localized spatial cues and temporal signals to out-of-image facts such as dates, events, or locations.
A general, model-agnostic agent framework that interleaves text/image search with a Set-of-Mark (SoM) zoom-in pathway, enabling region-seeded retrieval (zoom_in/image_search) and systematic analysis of how cropping and “thinking with images” affect current MLLMs.
We next motivate why such a benchmark is needed. Multimodal browsing benchmarks such as MMSearch (Jiang et al., 2024) primarily test an MLLM’s ability to coordinate text and images, but many of their questions admit relatively fixed workflows. A strong image search engine often retrieves pages whose surrounding text directly contains the answer, allowing unimodal LLMs—without visual perception—to perform competitively. In these cases, multimodality collapses to a narrow form of image–source cross-validation, and fine-grained visual reasoning plays only a minor role.
In contrast, recent text-only browsing tasks such as BrowseComp (Wei et al., 2025) emphasize persistence, multi-step evidence gathering, and complex search strategies. State-of-the-art MLLMs score below 2% with browsing tools, indicating substantial headroom and revealing a discrepancy: multimodal browsing benchmarks are far easier than their text-only counterparts, even though real-world multimodal tasks often require deeper reasoning.
Design principles for a challenging multimodal browsing benchmark.
Difficult web-based visual tasks typically exhibit three recurring challenges:
1.
Noisy or conflicting retrieval. Image search may surface near-duplicates, irrelevant scenes, or contradictory candidates; solving the task requires cross-source consistency checks and multimodal verification.
2.
Exhaustive, part-based visual reasoning. When holistic matches fail, agents must reason over subregions—cropping, zooming, and re-searching—to accumulate fragmented spatial cues.
3.
Long, tool-augmented reasoning chains. Realistic tasks require sustained coordination of visual tools (cropping, OCR), textual browsing, and multimodal evidence integration.
Benchmarks such as MMSearch do not consistently elicit these behaviors: many images contain a single dominant entity that quickly anchors a search, and the subsequent steps are primarily text-based.
MMSearch-Plus.
To construct tasks that inherently require fine-grained visual reasoning, MMSearch-Plus introduces Spatial–Temporal Extrapolation: we curate images containing multiple subtle signals—micro-text, uniform variations, broadcast overlays, layout signatures—that jointly imply facts not explicitly present in the prompt. Models must extract these cues, propagate them through iterative searches, and verify provenance amid retrieval noise. This better reflects real-world scenarios where neither the prompt nor the image alone suffices; for instance, determining the date of an esports broadcast or identifying a specific match from nearly identical sports frames. Finally, we develop a general agent framework with Set-of-Mark (SoM) to study how region-level visual grounding influences browsing agents. SoM provides human-verified marks and region-seeded queries, enabling “thinking with images” through structured zoom-in reasoning. Ablations quantify how cropping and localized visual extraction affect current MLLMs.
Together, MMSearch-Plus provides a high-difficulty, multimodal benchmark and analysis toolkit for evaluating the next generation of multimodal browsing agents. It differs from concurrent efforts such as BrowseComp-VL (Geng et al., 2025) and MM-BrowseComp (Li et al., 2025b) in both data sources and its agent framework, which emphasizes sustained fine-grained visual reasoning.
2 Method
2.1 Spatial-Temporal Extrapolation
Summary. We test an agent’s ability to infer facts that are not literally present in the pixels or the instant shown, requiring it to recover the broader event by combining partial visual cues with retrieval and reasoning across time and space. This emphasis discourages purely perceptual shortcutting and directly measures uncertainty-aware planning, precise event localization, and evidence-backed browsing.
A central challenge in BrowseComp-like benchmarks is the ballooning intermediate search space induced by soft, fuzzy constraints, which demands non-trivial cross-validation to identify the correct target. Rather than remixing text-only corpora, we construct multimodal problems that naturally expand this search space: from a handful of visual fragments tied to real-world events (in the spirit of GeoGuessr AB (2013)), the agent must hypothesize the underlying source event and verify it against retrieved evidence. Difficulty is modulated by the richness of visual and textual cues; even a single crop can widen the candidate set dramatically, eliciting iterative hypothesis–test–refine behavior and long trajectories that require robust context tracking.
Once the source event is pinned down, our questions probe metadata and support multi-hop reasoning. To elevate difficulty, we introduce Spatial-Temporal Extrapolation: instead of querying what is directly visible, we ask what is contextually implied but physically absent, compelling reasoning beyond the frame and moment to reconstruct the broader situation. Spatial extrapolation targets unseen entities (e.g., off-screen participants, individuals facing away, occluded signage), while temporal extrapolation probes events immediately before or after the depicted moment (e.g., lineups, the next play/goal, or the subsequent segment/episode). Solving these requires agents to (i) precisely localize the event (time, match, or episode), and (ii) retrieve and integrate wider contextual knowledge from diverse sources to produce an evidence-grounded answer.
2.2 Data Collection
Building on the Spatial–Temporal Extrapolation framework, we instantiated the data collection process as follows.
We began with a set of predefined categories and generated search keywords using a mix of LLM assistance and manual curation. Data were sourced from publicly accessible video-sharing platforms and open-access scholarly repositories (i.e., arXiv). Queries for consumer video content (e.g., Vlog) were issued to video platforms, while those in the Acad. category were issued to arXiv to retrieve relevant research papers (PDFs). All materials are used only for research. We also mask sensitive content (e.g., faces and vehicle license plates). In our pipeline, both videos and papers are treated as events.
For videos, human annotators manually selected keyframes suitable for constructing questions. The keyframes underwent anonymization for sensitive content. For papers, annotators extracted figures or tables, capturing them via screenshots. A screenshot was considered suitable if it (i) contained vague or noisy information that challenged perception, (ii) depicted less familiar entities requiring external retrieval to answer associated questions, and (iii) could not be searched with Google image search directly. This ensures that most, if not all, screenshots are not in the training dataset, and the essential information for answering the question is neither in the training corpora nor directly retrievable via a direct image search.
Filtering.
Although the collection focused on new or rarely seen events, we observed that certain closed-source models (e.g., GPT-4o, GPT-5, Gemini-2.5-Pro) could occasionally solve questions without external search. To counteract this, we employed an adversarial filtering procedure to increase difficulty. Specifically, annotators removed or obfuscated questions answerable without retrieval through the following methods:
Cross-validation: Annotators designed problems outside their own knowledge base, then verified solvability by testing on at least two closed-source MLLMs.
Image perturbation: Critical visual subregions were blurred or masked.
Iterative refinement: Questions that remained trivially solvable were discarded.
Dataset updates.
Some problems that originally required retrieval later became solvable without search because newer closed-source MLLMs began to recognize previously unfamiliar entities or figures. This temporal drift is unavoidable as models continue to incorporate more recent data. We therefore commit to regularly refreshing the benchmark to suppress such internal-knowledge shortcuts.
2.3 Dataset Statistics
| Statistic | Number |
| Total questions w/ images | 311 |
| Primary categories | 8 |
| Secondary categories | 43 |
| Unique images | 441 |
| Unique questions | 279 |
| Unique answers | 306 |
| Maximum question length | 42 |
| Maximum answer length | 20 |
| Average question length | 12.5 |
| Average answer length | 3.7 |
| Difficulty distribution | |
| - Easy | 94 (30.2%) |
| - Difficult | 217 (69.8%) |
Figure 3 depicts the distribution of MMSearch-Plus tasks across eight primary categories, which is approximately balanced. Each task includes at least one image, with additional images provided when a single view is insufficient. The benchmark primarily comprises short-answer questions, with answers averaging 3.7 words. In hindsight, we labeled the difficulty level of a task as Easy if it can be answered correctly without search or with only image search by o3 or Gemini-2.5-Pro. Otherwise, the task is labeled as Difficult (interchangeably Hard).
2.4 Benchmark Metric and Evaluation
We adopt accuracy as the primary evaluation metric. For each question, a set of acceptable answers is provided. We use LLM-as-a-judge to compare model predictions against this set, deeming an answer correct if it matches any acceptable variant. Validation through manual inspection showed complete agreement between human judgment and the assessment of GPT-4o.
2.5 Search Agent Framework
Overview.
We implement a model-agnostic web agent that interleaves text search, image search, and a zoom-in tool. Ranked results are obtained via SerpApi ; the surrounding framework (i.e., the system that includes the MLLM agent) performs lightweight post-processing to summarize pages, relate images to the query, and manage multi-round context. To reduce latency and token cost across repeated steps/evaluations, the framework caches image-search returns (ranked URLs/thumbnails) and their MLLM summaries per benchmark image. The framework maintains a threaded conversation state (tool calls, cropped views, summaries, and hypotheses) to support long-horizon, contextualized reasoning. It is compatible with sub-image cropping and sub-image search for fine-grained analysis.
Text and image search.
At any step, the agent may invoke text or image search. SerpAPI returns ranked candidates, from which the framework forwards a small set of top results to the model (images are summarized and cached; webpages are summarized). Retrieved webpages are summarized by an MLLM (Gemini in our runs) into two fields:
web_info: task-conditioned semantic summaries of page content that may support answering the current question;
related_info: evidence linking a result’s thumbnail/lead image(s) to the query image (e.g., matching signage, layouts, or micro-text).
These summaries are designed to compress interaction history for use within limited MLLM context windows; metadata such as URL/title is readily handled by SerpAPI. All retrieved and generated artifacts are integrated into the conversation state to enable contextualized evaluation across turns.
Zoom-in via Set-of-Mark (SoM).
To enable precise, provenance-aware inspection, we augment the agent with a Set-of-Mark (SoM) module (Yang et al., 2023). In this work, we refer to SoM as a list of human-verified bounding boxes per task image (examples in Fig. 9). In the first turn, we present both the raw image and an overlaid version with box outlines and indices, avoiding occlusion while allowing cross-reference.
3 Experiment
Evaluated Models.
We evaluate both closed-source and open-source MLLMs on our benchmark. Closed-source models include Gemini-2.5-Pro (hereafter Gemini) (Team and DeepMind, 2025), o3 (OpenAI, 2025b) and GPT-5 (OpenAI, 2025a), representing the current frontier of proprietary multimodal systems. On the open-source side, we focus on Qwen-2.5-VL-72B-Instruct (hereafter Qwen) (Bai et al., 2025). Expert human annotators were instructed to work with a Chrome browser for at least 10 minutes for each task, and capped at 20 minutes.
| Model / Search Mode | Avg | By Category | Difficulty | ||||||||
| Geo. | Sports | Acad. | Film/TV | Tech | Games | Vlog | Music | Easy | Hard | ||
| Human | |||||||||||
| Browser | 22.8 | 20.3 | 25.9 | 20.0 | 25.0 | 19.4 | 16.1 | 31.6 | 35.3 | 34.0 | 18.0 |
| Closed-source | |||||||||||
| o3 (2025-04-16) | |||||||||||
| Without Search | 15.1 | 31.2 | 14.8 | 6.0 | 17.5 | 13.9 | 3.2 | 5.3 | 11.8 | 50.0 | 0.0 |
| Image Search | 19.3 | 28.1 | 14.8 | 18.0 | 30.0 | 22.2 | 3.2 | 5.3 | 17.6 | 63.8 | 0.0 |
| Text Search | 37.0 | 43.8 | 35.2 | 48.0 | 30.0 | 44.4 | 16.1 | 31.6 | 29.4 | 50.0 | 31.3 |
| Full Rollout | 36.0 | 35.9 | 24.1 | 50.0 | 42.5 | 44.4 | 16.1 | 42.1 | 29.4 | 54.3 | 28.1 |
| Full Rollout + SoM | 37.6 | 45.3 | 29.6 | 46.0 | 45.0 | 45.7 | 16.1 | 26.3 | 29.4 | 62.8 | 26.9 |
| GPT-5 | |||||||||||
| Without Search | 10.3 | 21.9 | 7.4 | 4.0 | 7.5 | 8.3 | 0.0 | 5.3 | 15.8 | 27.7 | 2.8 |
| Image Search | 16.4 | 25.0 | 11.1 | 14.0 | 22.5 | 19.4 | 3.2 | 0.0 | 29.4 | 50.0 | 1.8 |
| Full Rollout + SoM | 35.4 | 35.9 | 27.8 | 48.0 | 50.0 | 41.7 | 6.5 | 36.8 | 23.5 | 56.4 | 26.3 |
| Gemini-2.5-Pro | |||||||||||
| Without Search | 10.6 | 15.6 | 11.1 | 6.0 | 12.5 | 13.9 | 0.0 | 15.8 | 5.9 | 35.1 | 0.0 |
| Image Search | 16.4 | 26.6 | 11.1 | 18.0 | 20.0 | 16.7 | 3.2 | 0.0 | 23.5 | 54.3 | 0.0 |
| Full Rollout | 23.8 | 39.1 | 14.8 | 12.0 | 27.5 | 33.3 | 6.5 | 26.3 | 29.4 | 46.8 | 13.8 |
| Full Rollout + SoM | 27.7 | 40.6 | 22.2 | 24.0 | 25.0 | 33.3 | 19.4 | 15.8 | 29.4 | 54.3 | 16.1 |
| Open-source | |||||||||||
| Qwen-2.5-VL-72B | |||||||||||
| Without Search | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Image Search | 13.5 | 20.3 | 7.4 | 18.0 | 17.5 | 11.1 | 3.2 | 0.0 | 23.5 | 41.5 | 1.4 |
| Full Rollout | 6.1 | 9.5 | 7.4 | 4.0 | 5.0 | 2.8 | 3.2 | 5.3 | 11.8 | 17.0 | 1.4 |
| Full Rollout + SoM | 7.1 | 10.9 | 3.7 | 4.0 | 10.0 | 5.6 | 6.5 | 5.3 | 11.8 | 18.1 | 2.3 |
Implementation Details.
To ensure a fair comparison, we standardize the search-and-reasoning pipeline across all models. We consider five search modes:
Without Search. The MLLM answers using only the question and image inputs.
Image Search. For each benchmark image, the top-10 image-search results are summarized by Gemini and cached. The MLLM then receives the question, the images, and these summaries as input. For multi-image questions, we append each image’s summaries to the model context.
Text Search. For text search, we retrieve the top-5 websites by default. The process is capped at 20 search rounds. In each text-search round, the model generates 3–5 refined queries
Full Rollout. The model runs our search-agent framework and may call search tools as needed. Image-search and text-search settings match those described before. The process is capped at 20 search rounds (combining text and image search) to balance efficiency and coverage. In each image-search round, it issues exactly one query. For Qwen, pilot studies showed that more than 10 rounds yield no further gains, so we cap its search at 10 rounds.
Full Rollout + SoM. Images overlayed with Set-of-Mark are provided to MLLMs. In addition to capabilities provided in Full Rollout, models can (a) call the zoom-in tool to examine a subregion of an image, (b) call the image search tool to search with the subregion. The maximum number of tool calls is 20.
4 Results
The results of MLLMs across search modes are presented in Table 1.
Baseline performance.
Under No Search on the original MMSearch+ benchmark, open-source Qwen scored 0.0%, confirming that MMSearch+ resisted parametric shortcuts. Closed-source models remained far from saturation (GPT-5 10.3, Gemini-2.5-Pro 10.6, o3 15.1). With Image Search (o3 +4.2, GPT-5 +6.1, Gemini +5.8; Qwen +13.5), models gained from coarse disambiguation, yet multi-hop composition remained the main blocker. With iterative image and text search, Full Rollout produced large gains: o3 36.0 (+20.9 over no-search) and Gemini 23.8 (+13.2), while Qwen regressed to 6.1 (-7.4 over image-search). This indicates that the external information retrieved from search tools is essential to answering the questions correctly.
Performance drop on easy split.
Notably, easy split accuracy did not improve from Image Search to Full Rollout. For o3, we conducted a targeted error analysis on this split: there are 23 “easy” questions for which the image-only setting is correct but the full-rollout setting is not, corresponding to a 7.4-point absolute accuracy loss on our 311-task benchmark. We randomly sampled 10 of these 23 cases and manually inspected the trajectories. In 9 out of 10, the full-rollout trajectory never invoked image search at all: the model apparently believed it had already understood the image well enough and instead relied solely on text search or prior knowledge, thereby missing critical fine-grained visual cues. In the remaining case, the model did perform image search and retrieved relevant sources, but made a reasoning error in the final step.
These observations indicate that the drop in Easy-split accuracy is not primarily driven by over-retrieval or exposure to distractors, but by underuse of image search in the more complex full-rollout setting: the model sometimes “skips” a necessary image-search step when it incorrectly assumes that additional retrieval is unnecessary. We therefore attribute the Easy-split gap mainly to:
1.
Label vs. retrieval difficulty. In MMSearch+, “easy” meant higher likelihood of being in-model or solvable by a direct image hop, not that multi-step retrieval was trivial.
2.
Tool-use decision errors. The dominant failure mode for o3 is a policy-level issue—deciding whether to invoke image search—rather than over-retrieval or inherent drawbacks of the full-rollout setting.
Overall, rollout helped where progressive narrowing was necessary (academic, tech, media queries), but on easy items performance was limited by miscalibration in tool choice: the model occasionally declined to use image search even when it was crucial for success.
Adding SoM (Full Rollout + SoM) gave consistent, targeted gains.
The improvements were o3 37.6 (+1.6), Gemini 27.7 (+3.9), Qwen 7.1 (+1.0) over Full Rollout. Set-of-Mark anchored reasoning on fine-grained visual cues (micro-text, logos, key entities) before expanding the search space.
Decoupling internal knowledge.
To disentangle the effect of models’ internal knowledge from their use of external tools, we further evaluate on MMSearch, a subset of 239 tasks that are unsolvable without search by any model we tested (Table 4). The relationship between performance on the full and lite benchmarks is visualized in Figure 4.
On MMSearch, the best-performing setting is full rollout with only text search (abbreviated as Text Search in the table) with the o3 model, achieving 31.4% accuracy. This suggests that, once questions solvable purely from internal memory are removed, current MLLMs still struggle to exploit fine-grained image search effectively. In many categories, we observe a drop in accuracy when moving from only text search to full rollout / full rollout + SoM. However, there remain categories where image search is clearly beneficial: for o3 in Sports, accuracy increases by 6.6 (full rollout) and 10.0 points (full rollout + SoM) over text-only rollout. Overall, most methods lie close to the diagonal in Figure 4, indicating that the qualitative trends observed on MMSearch+ are not driven primarily by parametric memorization, but by genuine tool-use and retrieval behavior.
5 Analysis
5.1 Observations
Closed-source models retained broad geographic knowledge.
They often recognized rare locales and supplied plausible context; geography tasks were comparatively strong. Without search tools, o3 achieved 31.2%, while human annotators with a standard browser only achieved 20.3%. On other categories like sports and Vlog, the human baseline usually outperformed the tested MLLMs. We name this phenomenon the different comparative advantages of humans and MLLMs.
o3 exhibited strong long-context coordination, while Gemini leveraged fewer search calls.
On the one hand, o3 sustained search rounds and reasoned over retrieved items without excessive redundancy, contributing to its rollout gains. On the other hand, Figure 5 (left) shows that Gemini always answered on or before Step 9. However, more tool calls did not correlate with trajectory correctness: Figure 15 shows that incorrect trajectories typically contained more search calls than correct ones.
Format correctness of closed-source models’ tool calls, and Qwen’s tool-use instability.
All the closed-source models consistently produced valid tool calls. For Qwen, we observed invalid image-search calls across of tasks, inflating retries without quality improvements.
More text search calls than image search calls, and zoom-in operations were not always followed by a sub-image search call.
In Figure 5, all models made more text-search calls than image-search calls. We analyze the contribution of the “zoom-in then image-search” tool-call pattern to the final accuracy. For Gemini-2.5-Pro, out of 39 newly answered correctly questions from full rollout to full rollout + SoM, there are 5 such cases in which the model zoomed in on an image and then performed an image search (12.8%). For o3, out of 42 newly answered correctly questions, there is 1 such case (2.4%). This indicates that Gemini utilized fine-grained image search more frequently, while o3 often zoomed in simply to view a region more clearly.
To empirically support this observation, we examine the Markov transition probabilities between tool types (Figure 12). In particular, Set-of-Mark allows cropping and sub-image search, so we measure how often a zoom action is immediately followed by an image search. The transition probability varies substantially across models: Gemini uses image search after zooming in 25.37% of cases (34/134), Qwen in 10.56% (15/142), and o3 in only 2.87% (18/627). These statistics are consistent with the accuracy gains above—Gemini tends to use zoomed-in regions as retrieval queries, whereas o3 seldom does, often preferring further zooms or text search instead.
Image search and text search benefit different task categories.
The accuracies of o3 with text search (up to 10 rounds) and with full rollout are similar (37.0 vs. 36.0). However, their category-level accuracy distributions differ substantially. For example, image search yields better performance in Film/TV, whereas additional text search rounds provide greater benefits in Geography.
Scene-text is predominantly processed as textual evidence rather than visual evidence.
Across tasks containing logos, signage, or other forms of scene text, we observed that MLLMs overwhelmingly chose to convert visual text into explicit textual queries rather than invoking image search on the corresponding subregions. By detecting scene text with DeepSeek-OCR and manually inspecting sampled trajectories, we found that both o3 and Gemini-2.5-Pro typically relied on their built-in OCR capabilities to extract strings and issue text-search calls. For o3, 8 out of 10 sampled trajectories used only text search, while Gemini behaved similarly in 7 of 10 cases. Direct image-search calls on SoM-highlighted text regions were rare and generally appeared only when the model seemed uncertain about the exact transcription or wished to leverage additional visual cues such as layout or font. These findings suggest that, within MMSearch-Plus, scene-text reasoning is mediated primarily through textual grounding rather than visual retrieval, highlighting that multimodal agents implicitly choose between two tool pathways, (a) OCR + text search versus (b) image search, based on confidence and perceived utility.
5.2 Error Analysis
We categorize failure cases using a manually curated taxonomy and report the distribution over categories in Figure 6. Labels are assigned by the authors (single dominant label per erroneous prediction) for Gemini-2.5-Pro under the full-rollout setting. Examples of incorrect task trajectories are provided in Figure 7.
High-frequency categories.
The two most prevalent failure modes are no relevant information found and hallucination, responsible for and of errors, respectively ( combined). The former typically reflects a retrieval failure: noisy or overly specific queries miss on-topic pages. The latter reflects vision-driven misattribution: the model latches onto an incorrect event from internal knowledge and either does not invoke search or discounts contradictory evidence in the retrieved pages.
Other observed categories.
We also observe failures from misinterpreting the multimodal context; hallucinations grounded in either images or parametric knowledge; refusals or multi-answer outputs when evidence is insufficient to disambiguate a single event; failing to invoke search (falling back on outdated priors); and errors when summarizing relevant webpages.
Qualitative takeaways and implications.
Two recurring patterns emerge: (i) there is no single “golden path”—agents can succeed via varied orders of search/crop/verify; and (ii) stronger multimodal internal knowledge narrows the practical search space, enabling targeted verification over broad exploration. While augmenting multimodal internal knowledge can reduce exploration overhead, end-to-end reliability still depends on robust verification and tighter cross-modal grounding.
6 Related Work
Classical vision–language benchmarks probe compositional and commonsense reasoning in images and text (Johnson et al., 2017; Krishna et al., 2016; Zellers et al., 2019; Mathew et al., 2021). Among these, GQA (Hudson and Manning, 2019) emphasizes structured compositional reasoning using real-world scenes, while OK-VQA (Marino et al., 2019) requires external knowledge beyond the visual signal. More recently, InfoSeek (Chen, 2023) frames VQA as an information-seeking task, but relies on offline retrieval corpora whose contents frequently overlap with pretraining data for modern MLLMs, weakening its ability to measure genuine retrieval behavior. LiveVQA (Fu et al., 2025) benchmarks live visual knowledge by deriving VQA instances from recent web content beyond models’ cutoff dates, and shows that multimodal search augmentation substantially boosts performance.
Beyond static images, recent benchmark suites extend to long videos and interleaved modalities (Wu et al., 2024; Li et al., 2024; Hu et al., 2025; Jeong et al., 2025). Concurrent work on web agents studies planning, browsing, and evidence gathering in open worlds (Krishna et al., 2024; Wei et al., 2025; Geng et al., 2025; Li et al., 2025a; Tao et al., 2025).
Multimodal web search benchmarks begin to couple vision with browsing but often emphasize entity matching or template-based grounding over fine-grained visual reasoning (Jiang et al., 2024; Wu et al., 2025). Closest to our setting, MM-BrowseComp targets image/video evidence encountered during search and reveals substantial headroom even for frontier models (Li et al., 2025b).
A complementary line promotes “thinking with images” via improved grounding or RL-style training to elicit image-first reasoning (Wu and Xie, 2024; Zheng et al., 2024; Fan et al., 2025; Zheng et al., 2025b; Su et al., 2025a; Xia et al., 2025).
We differ from prior benchmarks by targeting sparse, spatial–temporal visual cues that must be transformed into targeted queries and multi-step browsing plans. Crucially, tasks in MMSearch-Plus are designed to be extremely unlikely to be answerable from internal model knowledge alone, requiring models to (i) decide what visual regions matter, (ii) choose appropriate external tools (image vs. text search), and (iii) gather and verify evidence online. This provides a focused stress test for agentic multimodal browsing systems.
7 Conclusion
This work introduces MMSearch-Plus, a benchmark that restores the genuinely multimodal pressures of web browsing by requiring agents to extract weak, localized visual cues and propagate them through noisy retrieval to out-of-image facts. Our Spatial–Temporal Extrapolation curation procedure preserves the persistence and difficulty profile of text-only browsing while compelling fine-grained visual reasoning and provenance verification. Evaluating a range of open and closed MLLMs within a model-agnostic search agent framework reveals a substantial headroom: the strongest agent we tested (o3) achieved less than 40% accuracy, while a strong open-source model (Qwen-2.5-VL-72B-Instruct) achieved 13.5%. These results indicate that current systems struggle to (i) decide when and how to crop and reuse visual evidence, and (ii) maintain verifiable chains of evidence across long-horizon tool use. We hope MMSearch-Plus serves as a rigorous stress test and a common yardstick for progress on multimodal browsing.
Ethics Statement
Principles.
We follow the Code of Ethics: uphold scientific rigor, avoid harm, ensure transparency, foster fairness, respect intellectual labor, and protect privacy and confidentiality.
Data sources, consent, and licensing.
We curate only publicly accessible material from (i) consumer video platforms (keyword queries) and (ii) open-access repositories (e.g., arXiv). We comply with platform Terms of Service and licenses and preserve attributions (e.g., video URL, paper URL). Items are treated as “events” and used solely for research.
Privacy, minimization, and redaction.
We mask sensitive content (e.g., faces, license plates); we exclude minors and content revealing private spaces when feasible. Residual risks remain; we provide an appeals/takedown channel.
Fairness.
To counter skews in public platforms, we conducted manual sampling across predefined categories and regions, report measurable distributions, and forbid uses that target protected classes or amplify inequities.
Limitations.
While diverse, our data inherits biases from public web sources and the domains we sampled; some tasks still under-represent non-English or low-visual-density pages. Our evaluation focuses on general-purpose agents with standard tool access; specialized systems with richer UI control or domain knowledge may behave differently. Finally, we primarily target image–text browsing; videos and dynamic interfaces remain underexplored.
Reproducibility Statement
We have taken several steps to facilitate independent verification of our results. Implementation details, hyperparameters, and evaluation protocols are documented in Section 3. We uploaded a randomly selected subset (seed = 42, 30 task examples) as supplementary material due to the 100MB space constraint, and released an anonymized archive on how to use the dataset in https://anonymous.4open.science/r/MMSearch-Plus-D54E.
Appendix A Use of LLMs
The manuscript text was polished for fluency using Grammarly after the authors completed a full draft. The authors then manually reviewed all suggestions to ensure their appropriateness.
Appendix B Example from MMSearch-Plus
Appendix C Comparison with Previous Benchmarks
| Benchmark/ Dataset | MM | Search | Multi-hop (2) | Thinking w/ Image | Multi- Image | ||
| Text | Image | Multi- turn | |||||
| InfoSeek (Chen, 2023) | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ |
| FRAMES (Krishna et al., 2024) | ✗ | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ |
| MMSearch (Jiang et al., 2024) | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ |
| FactualVQA(Wu et al., 2025) | ✓ | ✗ | ✓ | ✗ | ✗ | ✗ | ✗ |
| ChineseSimpleVQA (Gu et al., 2025) | ✓ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ |
| BrowseComp (Wei et al., 2025) | ✗ | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ |
| BrowseComp-ZH (Zhou et al., 2025) | ✗ | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ |
| BrowseComp-VL (Geng et al., 2025) | ✓ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ |
| MM-BrowseComp (Li et al., 2025b) | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✗ |
| Ours | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
Appendix D Related Work
We present a more comprehensive literature review in the Appendix.
Multimodal Reasoning Benchmarks.
Classical vision-and-language benchmarks such as CLEVR (Johnson et al., 2017), VCR (Zellers et al., 2019), Visual Genome (Krishna et al., 2016), and InfographicVQA (Mathew et al., 2021) established core evaluations for compositional reasoning, commonsense inference, and visual question answering. Recent efforts extend these settings to longer contexts and interleaved modalities, including LongVideoBench (Wu et al., 2024), MVBench (Li et al., 2024), Video-MMMU (Hu et al., 2025), and VideoRAG (Jeong et al., 2025), which highlight long-context understanding and retrieval augmentation. Other work such as MIRB (Zhao et al., 2024), pixel-level VCR (Zhang et al., 2024), and ReasonMap (Feng et al., 2025) aim to connect perception with explicit reasoning. However, these datasets are primarily self-contained and do not require dynamic interaction with the open web or tool-mediated search.
Web Browsing and Search-Enhanced Agents.
A large body of work explores language models as web agents that plan, browse, and gather evidence. Earilier works like FRAMES (Krishna et al., 2024) focus on multi-hop QAs on Wikipedia-like data. More recently, benchmarks such as BrowseComp (Wei et al., 2025) and its Chinese variant BrowseComp-ZH (Zhou et al., 2025) test multi-step persistence in text-heavy web navigation, where even frontier models struggle. Open-source agents including WebSailor (Li et al., 2025a), WebShaper (Tao et al., 2025), and WebWatcher (BrowseComp-VL) (Geng et al., 2025) emphasize robustness, data synthesis, and limited multimodal inputs.
In parallel, reinforcement-learning–based frameworks push the boundaries of search adaptivity and depth. Search-R1 (Jin et al., 2025), R1-Searcher (Song et al., 2025), ReSearch (Chen et al., 2025), Pangu DeepDiver (Shi et al., 2025), and DeepResearcher (Zheng et al., 2025a) train persistence and decomposition skills through reward shaping. ASearcher (Gao et al., 2025) extends this direction with large-scale asynchronous reinforcement learning, decoupling exploration from evaluation to unlock long-horizon search over dozens of hops. Complementary agentic systems such as Open Deep Search (Alzubi et al., 2025), AutoAgent (Tang et al., 2025), and Search-o1 (Li et al., 2025c) further democratize and automate search-enhanced reasoning.
Multimodal Web Search and MM-BrowseComp.
Beyond text-only browsing, MMSearch (Jiang et al., 2024) and MMSearch-R1 (Wu et al., 2025) introduce image-grounded queries, but many tasks still reduce multimodality to entity matching rather than genuine visual reasoning. MM-BrowseComp (Li et al., 2025b) addresses this gap with a 224-question benchmark designed for multimodal browsing agents where key evidence may be embedded in images or videos encountered during search. The benchmark provides a verified checklist for fine-grained analysis of multimodal dependencies and reasoning paths, and reveals that even top models (e.g., OpenAI o3 with tools) achieve only accuracy, underscoring substantial headroom for native multimodal reasoning within browsing pipelines. Our setting is closely related to, and concurrent with, MM-BrowseComp, but differs in focusing on sparse visual cues that must be transformed into targeted queries rather than predominantly entity-level matching.
Thinking with Images.
A growing body of work argues that multimodality should go beyond tool orchestration to think with images. Methods such as V* (Wu and Xie, 2024) and instruction-guided masking (Zheng et al., 2024) improve visual grounding inside LLMs. RL-style training further enforces image-first reasoning: GRIT (Fan et al., 2025), DeepEyes (Zheng et al., 2025b), Pixel Reasoner (Su et al., 2025a), and Visionary-R1 (Xia et al., 2025) encourage explicit visual deliberation before action. Visual Agentic Reinforcement Fine-Tuning (Liu et al., 2025), OpenThinkIMG (Su et al., 2025b), and One-RL-to-See-Them-All (Ma et al., 2025) propose unified training frameworks for multimodal agents, while broader systems like PyVision (Zhao et al., 2025) and Thyme (Zhang et al., 2025) generalize the paradigm.
Positioning and Summary.
Taken together, web-browsing benchmarks stress persistence and decomposition in open-world settings (Wei et al., 2025; Li et al., 2025a; Tao et al., 2025; Geng et al., 2025), while multimodal benchmarks emphasize perception-heavy reasoning (Johnson et al., 2017; Krishna et al., 2016; Zellers et al., 2019; Mathew et al., 2021). Recent multimodal browsing datasets—including MM-BrowseComp (Li et al., 2025b)—begin to unite these threads but still leave open how sparse visual evidence should drive query formulation and multi-step planning. We formulate the Spatial-Temporal Extrapolation methodology to concretely design challenging VQAs for testing multimodal search abilities. We also stress the importance of grounding and cropping images to allow fine-grained multimodal reasoning.
Appendix E Detailed Dataset Statistics
| Category | Count | % |
| Geography | 64 | 20.6% |
| Sports | 54 | 17.4% |
| Academic Research | 50 | 16.1% |
| Film & TV | 40 | 12.9% |
| Technology | 36 | 11.6% |
| Video Games | 31 | 10.0% |
| Vlog | 19 | 6.1% |
| Music | 17 | 5.5% |
The distributions of primary categories across MMSearch-Plus are reported in Table 3.
Appendix F Set-of-Mark Implementation Details
Figure 9 shows a list of task images overlaid with Set-of-Mark bounding box annotations from MMSearch-Plus.
In our context, set-of-mark for a task image consists of bounding boxes of key visual entities. Each box has a unique index. In the first user turn, we present both the raw image and an overlaid version with box outlines and indices, avoiding occlusion while allowing cross-reference. The framework exposes:
zoom_in(index): return a cropped sub-image for localized reasoning;
image_search(index): launch an image search seeded by the cropped region (compatible with the standard image-search pathway).
We include SoM because baseline MLLMs were unreliable at proposing visually precise crops from scratch. Providing accurate marks reduces the burden on detection/localization and lets models focus on reading, matching, and reasoning. While automatic mark generation (e.g., strong open-vocabulary detectors) is conceivable, robust coverage is challenging due to long-tail entities common in our benchmark; therefore MMSearch-Plus uses human-curated boxes.
Appendix G Evaluation Details
We initially tested both GPT-4o and Gemini-2.5-Pro as judges and found their outputs consistent. The subsequent experiments used GPT-4o. To further support reproducibility, we optionally provide lightweight rule-based verification methods.
Appendix H Results on MMSearch
To eliminate the impact of model’s internal knowledge, we obtain MMSearch from MMSearch+ by removing all tasks from MMSearch+ answerable without search by any model we tested. The results are reported in Table 4 and visualized in Figure 10 and 4.
In MMSearch, the best performing setting is full rollout with only text search (abbreviated as Text Search in the table) with o3 model, achieving 31.4% accuracy. This suggests that the current MLLMs cannot cleverly utilize fine-grained image search. In many categories, we observed a drop in accuracy from only text search to full rollout/full rollout + SoM. However, in certain categories like Sports, there is a great benefit of image search: for o3, the accuracy increases by 6.6 (full rollout) and 10.0 points (full rollout + SoM).
| Model / Search Mode | Avg | By Category | |||||||
| Geo. | Sports | Acad. | Film/TV | Tech | Games | Vlog | Music | ||
| Human | |||||||||
| Browser | 18.8 | 11.6 | 26.7 | 17.9 | 33.3 | 22.0 | 20.7 | 6.7 | 26.7 |
| Closed-source | |||||||||
| o3 (2025-04-16) | |||||||||
| Without Search | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Image Search | 10.0 | 14.0 | 13.3 | 10.3 | 16.7 | 4.9 | 17.2 | 3.3 | 0.0 |
| Text Search | 31.4 | 44.2 | 16.7 | 41.0 | 25.0 | 29.3 | 37.9 | 16.7 | 26.7 |
| Full Rollout | 27.3 | 46.5 | 33.3 | 17.9 | 8.3 | 14.6 | 37.9 | 16.7 | 33.3 |
| Full Rollout + SoM | 30.1 | 44.2 | 36.7 | 30.8 | 16.7 | 24.4 | 34.5 | 16.7 | 20.0 |
| GPT-5 | |||||||||
| Without Search | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Image Search | 8.8 | 9.3 | 13.3 | 7.7 | 25.0 | 4.9 | 13.8 | 3.3 | 0.0 |
| Full Rollout + SoM | 28.5 | 46.5 | 36.7 | 23.1 | 8.3 | 24.4 | 34.5 | 6.7 | 33.3 |
| Gemini-2.5-Pro | |||||||||
| Without Search | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Image Search | 10.5 | 14.0 | 13.3 | 12.8 | 25.0 | 7.3 | 10.3 | 3.3 | 0.0 |
| Full Rollout | 17.2 | 4.7 | 23.3 | 28.2 | 25.0 | 12.2 | 27.6 | 6.7 | 20.0 |
| Full Rollout + SoM | 20.9 | 23.3 | 13.3 | 28.2 | 25.0 | 19.5 | 24.1 | 20.0 | 6.7 |
| Open-source | |||||||||
| Qwen-2.5-VL-72B | |||||||||
| Without Search | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Image Search | 8.4 | 14.0 | 10.0 | 7.7 | 25.0 | 2.4 | 10.3 | 3.3 | 0.0 |
| Full Rollout | 3.3 | 4.7 | 3.3 | 0.0 | 8.3 | 4.9 | 3.4 | 3.3 | 0.0 |
| Full Rollout + SoM | 4.2 | 2.3 | 6.7 | 2.6 | 8.3 | 4.9 | 3.4 | 6.7 | 0.0 |
Appendix I Agent Behavior Analysis
This appendix presents detailed visualizations of tool usage patterns across the three multimodal agents: Gemini, o3, and Qwen. All figures show episodes capped at 20 steps for readability.
I.1 Episode Raster Plots
Figure 11 shows the tool usage patterns for the 40 longest episodes from each agent. Each row represents one episode, sorted by length (longest at top). Columns represent sequential steps, with colors indicating tool types: text search, image search, zoom-in, and answer. Dark blue strips represent an MLLM did not fully utilize 20 rounds and stopped early.
I.2 Tool Transition Matrices
Figure 12 shows the Markov transition probabilities between tool types.
In particular, Set-of-Mark allows image cropping and sub-image searching. Therefore, we investigate whether many of the cropped images were searched in the immediate next round, i.e., whether is large.
Gemini: 25.37% (34 out of 134 ZOOM actions)
Qwen: 10.56% (15 out of 142 ZOOM actions)
o3: 2.87% (18 out of 627 ZOOM actions)
I.3 Zoom-In Behavior Analysis
Figures 13 and 14 analyze zoom-in behavior patterns across agents. o3 shows the most aggressive zooming behavior, while Qwen and Gemini are more conservative.
Appendix J Rollout Trajectory Statistics
Figure 15 shows the distribution of number of search calls for correct and incorrect task trajectories (left) and the relation of the number of words in a model’s trajectory (excluding the words for tool calls) vs. the number of search calls (right) under the Full Rollout setting.
Appendix K Good Cases
Appendix L Prompts